### v1

# Pluriclosed metrics on compact semisimple Lie groups

Abstract.
Given a compact semisimple Lie group $G$ and a maximal torus $T\subset G$, we give an explicit description of all left and $\operatorname{Ad}(T)$-invariant pluriclosed Hermitian structures on $G$ in terms of the corresponding root system. They depend on $12d+$ 1 parameters in the irreducible case, where $\dim{T}=2d$.
This research was partially supported by three grants from, respectively, CONICET, Univ. Nac. de Córdoba and Foncyt (Argentina)
1. Introduction
A Hermitian metric $g$ on a complex manifold $(M,J)$ is said to be pluriclosed (or SKT, short for strong Kähler with torsion) if the Kähler form $\omega=g(J\cdot,\cdot)$ satisfies that $dd^{c}\omega=0$ (we refer to the recent surveys [F, FG2] for further information). Being $dd^{c}\omega$ a $4$-form, the condition turns out to be very strong in high dimensions, specially in the homogeneous case, making the search for examples a hard task (see [AL, AN, FP, FS] for structure and classification results on solvmanifolds). The primal compact examples are bi-invariant metrics on any even dimensional compact semisimple Lie group $G$ endowed with any left-invariant complex structure. As known, these are precisely Bismut flat manifolds with finite fundamental group (see [WYZ]). It was proved in [FGV] that a compact homogeneous complex manifold with finite fundamental group (i.e., a C-space) admits an invariant pluriclosed metric only if it is a Lie group $G$ (modulo a flag manifold, i.e., a Kähler homogeneous manifold).
Any left-invariant complex structure $J$ on a compact semisimple Lie group $G$ is determined by the choices of a maximal torus $T\subset G$ and a set of positive roots $\Delta^{+}$ (see [S, W, Pi]), they are often called Samelson complex structures. Indeed, $J$ leaves invariant the $\operatorname{B}_{\mathfrak{g}}$-orthogonal decomposition $\mathfrak{g}=\mathfrak{t}\oplus\mathfrak{q}$ of the Lie algebra $\mathfrak{g}$ of $G$, where $\operatorname{B}_{\mathfrak{g}}$ is the Killing form of $\mathfrak{g}$, $J_{\mathfrak{t}}$ is any linear map such that $J_{\mathfrak{t}}^{2}=-I$ and on the complexification $\mathfrak{q}^{c}=\bigoplus\limits_{\alpha\in\Delta}{\mathbb{C}}E_{\alpha}$, the complex linear map determined by $J_{\mathfrak{q}}$ on $\mathfrak{q}^{c}$ is given by $J_{\mathfrak{q}}E_{\alpha}=\pm\mathrm{i}E_{\alpha}$ for all $\alpha\in\pm\Delta^{+}$. The space of all left-invariant complex structures on $G$ therefore depends on $2d^{2}$ parameters up to biholomorphism if $\dim{T}=2d$. Any left and $\operatorname{Ad}(T)$-invariant metric which is compatible with $J=(J_{\mathfrak{t}},J_{\mathfrak{q}})$ is of the form $g=g_{\mathfrak{t}}+g_{\mathfrak{q}}$, where $g_{\mathfrak{t}}$ is any inner product on $\mathfrak{t}$ compatible with $J_{\mathfrak{t}}$ and $g_{\mathfrak{q}}=(x_{\alpha})_{\alpha\in\Delta^{+}}$ is given by
$g_{\mathfrak{q}}=\sum_{\alpha\in\Delta^{+}}x_{\alpha}(-\operatorname{B}_{\mathfrak{g}})|_{\mathfrak{q}_{\alpha}},\qquad x_{\alpha}>0,\quad\mathfrak{q}_{\alpha}:=({\mathbb{C}}E_{\alpha}\oplus{\mathbb{C}}E_{-\alpha})\cap\mathfrak{g},$
producing a space of Hermitian metrics that depends on $d^{2}+|\Delta^{+}|$ parameters.
The following strong results on left-invariant pluriclosed metrics on $G$ were obtained by Fino and Grantcharov in [FG1]:
  1. (a)
    If $(G,J)$ admits a pluriclosed left-invariant metric, then it also admits a left and $\operatorname{Ad}(T)$-invariant pluriclosed metric $g$ for the maximal torus $T$ of $G$ attached to $J$.
  2. (b)
    For any left and $\operatorname{Ad}(T)$-invariant pluriclosed metric $g$, there exists a bi-invariant metric $g_{b}$ on $G$ such that $g_{\mathfrak{t}}=g_{b}|_{\mathfrak{t}}$. The proof is based on cohomological properties of the torus fibration $G\rightarrow G/T$ over the full flag manifold $G/T$.
  3. (c)
    A $5$-parametric space of pluriclosed metrics on $G=\mathrm{SO}(9)$ is given for certain complex structure $J$.
  4. (d)
    The only left and $\operatorname{Ad}(T)$-invariant pluriclosed metrics which are also CYT (Calabi-Yau with torsion, i.e., the Bismut Ricci form vanishes) are bi-invariant metrics.
As far as we know, beyond $\mathrm{SU}(2)\times\mathrm{SU}(2)$, $\mathrm{SU}(3)$ and $\mathrm{SO}(9)$, the only other known examples of pluriclosed left-invariant metrics (non bi-invariant) on compact semisimple Lie groups were exhibited in [P] for $G=\mathrm{G}_{2}$. The pluriclosed flow in this setting was studied in [B], obtaining as an application the non-existence of pluriclosed and CYT metrics beyond the bi-invariant (or Bismut flat) ones.
In this paper, after giving a formula for $dd^{c}\omega$ in terms of roots for any left and $\operatorname{Ad}(T)$-invariant Hermitian structure on $G$ (see Lemma 3.1), we provide an alternative algebraic proof for the result (b) above (see Proposition 3.4) and then we prove the following complete classification (see Theorem 3.5 for a more detailed statement). We fix a left-invariant complex structure $J$ attached to a maximal torus $T$ and consider the decomposition $\mathfrak{t}=\mathfrak{t}_{1}\oplus\dots\oplus\mathfrak{t}_{s}$, where the $\mathfrak{t}_{i}$ are the corresponding maximal tori of the simple factors of $G$.
Theorem 1.1.
Given a compact semisimple Lie group $G$ of rank $2d$ endowed with $J$, any left and $\operatorname{Ad}(T)$-invariant pluriclosed metric is given as follows: $g_{\mathfrak{t}}(\mathfrak{t}_{i},\mathfrak{t}_{j})=0$ for all $i\neq j$ and on each simple factor, up to scaling, $g_{\mathfrak{t}}=-\operatorname{B}_{\mathfrak{g}}|_{\mathfrak{t}}$ and $g_{\mathfrak{q}}=(x_{\alpha})_{\alpha\in\Delta^{+}}$, where
$x_{\alpha}:=\sum_{i=1}^{2d}k_{i}x_{i}-\left(\sum_{i=1}^{2d}k_{i}-1\right),\qquad\forall\alpha=\sum_{i=1}^{2d}k_{i}\alpha_{i}\in\Delta^{+},\quad k_{i}\in{\mathbb{Z}}_{\geq 0},$
$\Pi=\{\alpha_{1},\dots,\alpha_{2d}\}\subset\Delta^{+}$ is the set of simple roots and $x_{i}:=x_{\alpha_{i}}$ for $i=1,\dots,2d$ (equivalently, $x_{\alpha+\beta}=x_{\alpha}+x_{\beta}-1$ for all $\alpha,\beta,\alpha+\beta\in\Delta^{+}$).
Each of these metrics is compatible with a $d(d-1)$-parametric space of left-invariant complex structures on $G$. According to the theorem, for a fixed $J$, the space of all left and $\operatorname{Ad}(T)$-invariant pluriclosed metrics compatible with $J$ depends on $r(2d+1)$ parameters and contains an $r$-parametric space of bi-invariant metrics, where $r$ is the number of irreducible factors of $(G,J)$ as a complex manifold. In particular, any irreducible $J$ admits a unique compatible bi-invariant metric up to scaling. Note that $J$ can be irreducible on a non-simple $G$ (see Example 2.1).
Intriguingly enough, the last line in the theorem should be compared with the Kähler condition for the Hermitian structure $(J_{\mathfrak{q}},g_{\mathfrak{q}})$ on the flag manifold $G/T$, given by $x_{\alpha+\beta}=x_{\alpha}+x_{\beta}$ for all $\alpha,\beta,\alpha+\beta\in\Delta^{+}$, whose solutions are precisely
$x_{\alpha}:=\sum\limits_{i=1}^{2d}k_{i}x_{i},\qquad\forall\alpha=\sum\limits_{i=1}^{2d}k_{i}\alpha_{i}\in\Delta^{+}.$
We do not know whether there may be some hidden geometric reason for this behavior.
Potential applications of Theorem 1.1 include an alternative algebraic proof for result (d) above and the study of the pluriclosed flow as an ODE system for $(x_{1}(t),\dots,x_{2d}(t))$, which will be the object of further research.
Acknowledgements. We thank Beatrice Brienza for fruitful conversations.
2. Hermitian geometry of Lie groups
Let $G$ be a connected and compact semisimple Lie group of even dimension. The rank of $G$ is therefore also even, say $\operatorname{rk}(G)=2d$. We refer to §4 for the definitions and notations concerning the root system of $G$.
It was proved by Pittie [Pi] that any left-invariant complex structure on $G$ is one of the found independently by Samelson [S] and Wang [W], which are defined as follows. Given a maximal torus $T\subset G$ and an ordering $\Delta^{+}$ of roots, we consider $J:\mathfrak{g}\rightarrow\mathfrak{g}$ given by: the $\operatorname{B}_{\mathfrak{g}}$-orthogonal decomposition $\mathfrak{g}=\mathfrak{t}\oplus\mathfrak{q}$ is $J$-invariant, $J_{\mathfrak{t}}:=J|_{\mathfrak{t}}$ is any linear map such that $J_{\mathfrak{t}}^{2}=-I$ and if
$\mathfrak{g}^{c}=\mathfrak{t}^{c}\oplus\mathfrak{q}^{c},\qquad\mathfrak{q}^{c}=\bigoplus_{\alpha\in\Delta}{\mathbb{C}}E_{\alpha},$
then the complex linear map determined by $J_{\mathfrak{q}}:=J|_{\mathfrak{q}}$ on $\mathfrak{q}^{c}$ is given by
$J_{\mathfrak{q}}E_{\alpha}=\mathrm{i}E_{\alpha},\quad\forall\alpha\in\Delta^{+},\qquad J_{\mathfrak{q}}E_{\alpha}=-\mathrm{i}E_{\alpha},\quad\forall\alpha\in\Delta^{-}:=-\Delta^{+}.$
The biholomorphism class of the complex manifold $(G,J)$, $J=(J_{\mathfrak{t}},J_{\mathfrak{q}})$ does not depend on the choice of $T$ nor of $\Delta^{+}$, so we can fix $J_{\mathfrak{q}}$. Two different $J_{\mathfrak{t}}$’s produce biholomorphic complex structures if and only if they belong to the same orbit of certain finite subgroup of $\operatorname{Aut}(G)$ (see [Pi]). This implies that the space $\mathcal{C}^{G}$ of all left-invariant complex structures on $G$ depends on $2d^{2}$ parameters up to biholomorphism (recall that $\dim{T}=2d$).
The left and $\operatorname{Ad}(T)$-invariant metrics which are compatible with $J=(J_{\mathfrak{t}},J_{\mathfrak{q}})$ are all of the form $g=g_{\mathfrak{t}}+g_{\mathfrak{q}}$, where $g_{\mathfrak{t}}$ is any inner product on $\mathfrak{t}$ compatible with $J_{\mathfrak{t}}$ and $g_{\mathfrak{q}}=(x_{\alpha})_{\alpha\in\Delta}$ is given by
$g_{\mathfrak{q}}=\sum_{\alpha\in\Delta^{+}}x_{\alpha}(-\operatorname{B}_{\mathfrak{g}})|_{\mathfrak{q}_{\alpha}},\qquad\mathfrak{q}_{\alpha}:=({\mathbb{C}}E_{\alpha}\oplus{\mathbb{C}}E_{-\alpha})\cap\mathfrak{g}.$
In particular, $x_{\alpha}>0$ and we set $x_{-\alpha}:=x_{\alpha}$ for any $\alpha\in\Delta^{+}$. The corresponding symmetric ${\mathbb{C}}$-bilinear form on $\mathfrak{g}^{c}$, also denoted by $g$, satisfies that
$g(E_{\alpha},E_{-\alpha})=-x_{\alpha},\qquad\forall\alpha\in\Delta,$
and $g(E_{\alpha},E_{\beta})=0$ whenever $\alpha+\beta\neq 0$. The Killing metric $g_{\operatorname{B}}:=-\operatorname{B}_{\mathfrak{g}}$ has $x_{\alpha}=1$ for any $\alpha\in\Delta^{+}$ and a bi-invariant metric $g_{b}=z_{1}(-\operatorname{B}_{\mathfrak{g}_{1}})+\dots+z_{s}(-\operatorname{B}_{\mathfrak{g}_{s}})$ corresponds to $x_{\alpha}=z_{i}$ if and only if $\alpha\in\Delta^{+}_{i}$, where $G=G_{1}\times\dots\times G_{s}$ is the decomposition in simple factors and $\Delta_{i}$ is the root system of $G_{i}$.
We note that $(G,J)$ is irreducible as a complex manifold if and only if $J_{\mathfrak{t}}$ does not leave any subspace $\mathfrak{t}_{i_{1}}\oplus\dots\oplus\mathfrak{t}_{i_{k}}$ with $0<k<s$ invariant, where $\mathfrak{t}_{i}$ is the maximal torus of $\mathfrak{g}_{i}$. In that case, there exists a unique bi-invariant metric compatible with $J$ (up to scaling), which is not necessarily the Killing metric as the following example shows.
Example 2.1.
Consider $G=G_{1}\times G_{2}$ such that $G_{1},G_{2}$ are simple, $\operatorname{rk}(G_{1})=\operatorname{rk}(G_{2})$ and relative to the decomposition $\mathfrak{t}=\mathfrak{t}_{1}\oplus\mathfrak{t}_{2}$,
$J_{\mathfrak{t}}=\left[\begin{matrix}0&-\tfrac{1}{b}I\\ bI&0\end{matrix}\right],\qquad b\neq 0.$
Then $J$ is irreducible and a bi-invariant metric $g_{b}=z_{1}(-\operatorname{B}_{\mathfrak{g}_{1}})+z_{2}(-\operatorname{B}_{\mathfrak{g}_{s}})$ is compatible with $J$ if and only if $\tfrac{z_{1}}{z_{2}}=b$.
We consider the spaces $\mathcal{M}^{T}$, $\mathcal{C}^{T}$, $\mathcal{H}^{T}_{J}$, $\mathcal{C}^{T}_{g}$ and $\mathcal{H}^{T}$ of all left and $\operatorname{Ad}(T)$-invariant metrics, complex structures, metrics compatible with a fixed $J\in\mathcal{C}^{T}$, complex structures compatible with a fixed $g\in\mathcal{M}^{T}$ and Hermitian structures, respectively. They are all manifolds and it is easy to check that their dimensions are respectively given by
$\begin{array}[]{c}\dim{\mathcal{M}^{T}}=\tfrac{d(d+1)}{2}+|\Delta^{+}|,\qquad\dim{\mathcal{C}^{T}}=2d^{2},\\ \\ \dim{\mathcal{H}^{T}_{J}}=d^{2}+|\Delta^{+}|,\qquad\dim{\mathcal{C}^{T}_{g}}=d(d-1),\qquad\dim{\mathcal{H}^{T}}=3d^{2}+|\Delta^{+}|.\end{array}$
The Kähler form $\omega=g(J\cdot,\cdot)$ of a Hermitian structure $(J,g)$ on $G$ as above is given by
$\omega=\omega_{\mathfrak{t}}+\omega_{\mathfrak{q}},\qquad\mbox{where}\quad\omega_{\mathfrak{t}}=g_{\mathfrak{t}}(J_{\mathfrak{t}}\cdot,\cdot),\quad\omega_{\mathfrak{q}}(E_{\alpha},E_{-\alpha})=-\mathrm{i}\epsilon_{\alpha}x_{\alpha},\quad\forall\alpha\in\Delta,$
where $\epsilon_{\alpha}:=\pm 1$ for $\alpha\in\Delta^{\pm}$. Recall that if we consider $H_{\alpha}:=[E_{\alpha},E_{-\alpha}]\in\mathfrak{t}^{c}$, then the set $\{H_{\alpha}:\alpha\in\Delta^{+}\}$ generates $\mathfrak{t}^{c}$.
Lemma 2.2.
For any left and $\operatorname{Ad}(T)$-invariant Hermitian structure $(J,g)$ on a compact semisimple Lie group $G$, the only possibly nonzero components of the $3$-forms $d\omega$ and $d^{c}\omega:=-d\omega(J\cdot,J\cdot,J\cdot)$ are given by
$d\omega(E_{\alpha},E_{\beta},E_{\gamma})=N_{\alpha,\beta}(y_{\alpha}+y_{\beta}+y_{\gamma}),\qquad d^{c}\omega(E_{\alpha},E_{\beta},E_{\gamma})=\mathrm{i}\epsilon_{\alpha}\epsilon_{\beta}\epsilon_{\gamma}N_{\alpha,\beta}(y_{\alpha}+y_{\beta}+y_{\gamma}),$
whenever $\alpha+\beta+\gamma=0$, where $y_{\alpha}=-\mathrm{i}\epsilon_{\alpha}x_{\alpha}$ and $[E_{\alpha},E_{\beta}]=N_{\alpha,\beta}E_{\alpha+\beta}$ (see §4), and
$d\omega(A,E_{\alpha},E_{-\alpha})=\omega_{\mathfrak{t}}(A,H_{\alpha})=g_{\mathfrak{t}}(J_{\mathfrak{t}}A,H_{\alpha}),\qquad d^{c}\omega(A,E_{\alpha},E_{-\alpha})=g_{\mathfrak{t}}(A,H_{\alpha}),$
for all $A\in\mathfrak{t}^{c}$.
Remark 2.3.
In particular, $d\omega$ is always nonzero, in accordance with the well-known fact that $(J,g)$ is never Kähler. It also follows that the Hermitian structure $(J_{\mathfrak{q}},g_{\mathfrak{q}})$ on the flag manifold $G/T$ is Kähler if and only if $x_{\alpha+\beta}=x_{\alpha}+x_{\beta}$ for all $\alpha,\beta,\alpha+\beta\in\Delta^{+}$.
Proof.
For any $\alpha,\beta,\gamma\in\Delta$,
$d\omega(E_{\alpha},E_{\beta},E_{\gamma})$ $=-N_{\alpha,\beta}\omega_{\mathfrak{q}}(E_{\alpha+\beta},E_{\gamma})+N_{\alpha,\gamma}\omega_{\mathfrak{q}}(E_{\alpha+\gamma},E_{\beta})-N_{\beta,\gamma}\omega_{\mathfrak{q}}(E_{\beta+\gamma},E_{\alpha})$
$=N_{\alpha,\beta}y_{\gamma}-N_{\alpha,\gamma}y_{\beta}+N_{\beta,\gamma}y_{\alpha}=N_{\alpha,\beta}(y_{\alpha}+y_{\beta}+y_{\gamma}),$
if $\alpha+\beta+\gamma=0$ and zero otherwise. On the other hand, if $A\in\mathfrak{t}^{c}$, then
$d\omega(A,E_{\alpha},E_{\beta})$ $=-\alpha(A)\omega_{\mathfrak{q}}(E_{\alpha},E_{\beta})+\beta(A)\omega_{\mathfrak{q}}(E_{\beta},E_{\alpha})-\omega([E_{\alpha},E_{\beta}],A)$
=$=\left\{\begin{array}[]{l}0,\qquad\alpha+\beta\neq 0,\\ \omega_{\mathfrak{t}}(A,H_{\alpha}),\qquad\alpha+\beta=0.\end{array}\right.$
The only other possibilities, say $d\omega(A,B,E_{\alpha})$ and $d\omega(A,B,C)$ for $B,C\in\mathfrak{t}^{c}$, are easily seen to vanish, concluding the proof. ∎
3. Pluriclosed structures
A Hermitian structure $(J,g)$ is called pluriclosed if $dd^{c}\omega=0$ (or equivalently, ¯$\partial\overline{\partial}\omega=0$). We start this section by computing the $4$-form $dd^{c}\omega$ in our setting.
Lemma 3.1.
For any left and $\operatorname{Ad}(T)$-invariant Hermitian structure $(J,g)$ on a compact semisimple Lie group $G$, the only possibly nonzero components of the $4$-form $dd^{c}\omega$ are given by
$dd^{c}\omega(E_{\alpha},E_{-\alpha},E_{\beta},E_{-\beta})=$ $-2g_{\mathfrak{t}}(H_{\alpha},H_{\beta})-2N_{\alpha,\beta}^{2}(x_{\alpha+\beta}-x_{\alpha}-x_{\beta})$
$-2\epsilon_{\alpha-\beta}N_{\alpha,-\beta}^{2}(\epsilon_{\alpha-\beta}x_{\alpha-\beta}-x_{\alpha}+x_{\beta}),\qquad\forall\alpha,\beta\in\Delta^{+},\alpha\neq\beta,$
and if $\alpha+\beta+\gamma+\delta=0$ and all pairs add nonzero, then for all $\alpha,\beta\in\Delta^{+}$, $\gamma,\delta\in\Delta^{-}$,
$dd^{c}\omega(E_{\alpha},E_{\beta},E_{\gamma},E_{\delta})=$ $N_{\alpha,\beta}N_{\gamma,\delta}(x_{\alpha}+x_{\beta}+x_{\gamma}+x_{\delta}-2x_{\alpha+\beta})$
$-\epsilon_{\alpha+\gamma}N_{\alpha,\gamma}N_{\beta,\delta}(-x_{\alpha}+x_{\beta}+x_{\gamma}-x_{\delta}+2\epsilon_{\alpha+\gamma}x_{\alpha+\gamma})$
$+\epsilon_{\alpha+\delta}N_{\alpha,\delta}N_{\beta,\gamma}(-x_{\alpha}+x_{\beta}-x_{\gamma}+x_{\delta}+2\epsilon_{\alpha+\delta}x_{\alpha+\delta}).$
Remark 3.2.
Using the properties of the numbers $N_{\alpha,\beta}$ given in §4, one can easily check that $g_{\operatorname{B}}$ is therefore always pluriclosed. However, recall that there is a large class of left-invariant complex structures on $G$ which are not compatible with any bi-invariant metric.
Proof.
We use the formula for $d^{c}\omega$ given in Lemma 2.2. If $\alpha+\beta+\gamma=0$, then for any $A\in\mathfrak{t}^{c}$,
$dd^{c}\omega(A,E_{\alpha},E_{\beta},E_{\gamma})$
$=$ $-\alpha(A)d^{c}\omega(E_{\alpha},E_{\beta},E_{\gamma})+\beta(A)d^{c}\omega(E_{\beta},E_{\alpha},E_{\gamma})-\gamma(A)d^{c}\omega(E_{\gamma},E_{\alpha},E_{\beta})$
$-N_{\alpha,\beta}d^{c}\omega(E_{\alpha+\beta},A,E_{\gamma})+N_{\alpha,\gamma}d^{c}\omega(E_{\alpha+\gamma},A,E_{\beta})-N_{\beta,\gamma}d^{c}\omega(E_{\beta+\gamma},A,E_{\alpha})$
$=$ $-(\alpha+\beta+\gamma)(A)d^{c}\omega(E_{\alpha},E_{\beta},E_{\gamma})$
$-N_{\alpha,\beta}d^{c}\omega(A,E_{\gamma},E_{\alpha+\beta})+N_{\alpha,\gamma}d^{c}\omega(A,E_{\beta},E_{\alpha+\gamma})-N_{\beta,\gamma}d^{c}\omega(A,E_{\alpha},E_{\beta+\gamma})$
$=$ $-N_{\alpha,\beta}g_{\mathfrak{t}}(A,H_{\gamma})+N_{\alpha,\gamma}g_{\mathfrak{t}}(A,H_{\beta})-N_{\beta,\gamma}g_{\mathfrak{t}}(A,H_{\alpha})$
$=$ $-N_{\alpha,\beta}g_{\mathfrak{t}}(A,H_{\alpha}+H_{\beta}+H_{\gamma})=-N_{\alpha,\beta}g_{\mathfrak{t}}(A,H_{\alpha+\beta+\gamma})=0.$
For $A,B\in\mathfrak{t}^{c}$ and $\alpha+\beta=0$,
$dd^{c}\omega(A,B,E_{\alpha},E_{\beta})=$ $\alpha(A)d^{c}\omega(E_{\alpha},B,E_{\beta})-\beta(A)d^{c}\omega(E_{\beta},B,E_{\alpha})-\alpha(B)d^{c}\omega(E_{\alpha},A,E_{\beta})$
$+\beta(B)d^{c}\omega(E_{\beta},A,E_{\alpha})-N_{\alpha,\beta}d^{c}\omega(E_{\alpha+\beta},A,B)$
$=$ $-(\alpha+\beta)(A)d^{c}\omega(B,E_{\alpha},E_{\beta})+(\alpha+\beta)(A)d^{c}\omega(A,E_{\alpha},E_{\beta})=0.$
If $\alpha+\beta+\gamma+\delta=0$ and all pairs add nonzero, then
$dd^{c}\omega(E_{\alpha},E_{\beta},E_{\gamma},E_{\delta})$
$=$ $-N_{\alpha,\beta}d^{c}\omega(E_{\alpha+\beta},E_{\gamma},E_{\delta})-N_{\gamma,\delta}d^{c}\omega(E_{\gamma+\delta},E_{\alpha},E_{\beta})$
$+N_{\alpha,\gamma}d^{c}\omega(E_{\alpha+\gamma},E_{\beta},E_{\delta})+N_{\beta,\delta}d^{c}\omega(E_{\beta+\delta},E_{\alpha},E_{\gamma})$
$-N_{\alpha,\delta}d^{c}\omega(E_{\alpha+\delta},E_{\beta},E_{\gamma})-N_{\beta,\gamma}d^{c}\omega(E_{\beta+\gamma},E_{\alpha},E_{\delta})$
$=$ $-N_{\alpha,\beta}\mathrm{i}\epsilon_{\alpha+\beta}\epsilon_{\gamma}\epsilon_{\delta}N_{\gamma,\delta}(y_{\alpha+\beta}+y_{\gamma}+y_{\delta})-N_{\gamma,\delta}\mathrm{i}\epsilon_{\gamma+\delta}\epsilon_{\alpha}\epsilon_{\beta}N_{\alpha,\beta}(y_{\gamma+\delta}+y_{\alpha}+y_{\beta})$
$+N_{\alpha,\gamma}\mathrm{i}\epsilon_{\alpha+\gamma}\epsilon_{\beta}\epsilon_{\delta}N_{\beta,\delta}(y_{\alpha+\gamma}+y_{\beta}+y_{\delta})+N_{\beta,\delta}\mathrm{i}\epsilon_{\beta+\delta}\epsilon_{\alpha}\epsilon_{\gamma}N_{\alpha,\gamma}(y_{\beta+\delta}+y_{\alpha}+y_{\gamma})$
$-N_{\alpha,\delta}\mathrm{i}\epsilon_{\alpha+\delta}\epsilon_{\beta}\epsilon_{\gamma}N_{\beta,\gamma}(y_{\alpha+\delta}+y_{\beta}+y_{\gamma})-N_{\beta,\gamma}\mathrm{i}\epsilon_{\beta+\gamma}\epsilon_{\alpha}\epsilon_{\delta}N_{\alpha,\delta}(y_{\beta+\gamma}+y_{\alpha}+y_{\delta}).$
In particular, for $\alpha,\beta,\gamma\in\Delta^{+}$, $\delta\in\Delta^{-}$, we obtain
$dd^{c}\omega(E_{\alpha},E_{\beta},E_{\gamma},E_{\delta})$
$=$ $N_{\alpha,\beta}N_{\gamma,\delta}(x_{\alpha+\beta}+x_{\gamma}-x_{\delta})+N_{\gamma,\delta}N_{\alpha,\beta}(-x_{\gamma+\delta}+x_{\alpha}+x_{\beta})$
$-N_{\alpha,\gamma}N_{\beta,\delta}(x_{\alpha+\gamma}+x_{\beta}-x_{\delta})-N_{\beta,\delta}N_{\alpha,\gamma}(-x_{\beta+\delta}+x_{\alpha}+x_{\gamma})$
$+N_{\alpha,\delta}N_{\beta,\gamma}(-x_{\alpha+\delta}+x_{\beta}+x_{\gamma})+N_{\beta,\gamma}N_{\alpha,\delta}(x_{\beta+\gamma}+x_{\alpha}-x_{\delta})$
$=$ $(N_{\alpha,\beta}N_{\gamma,\delta}-N_{\alpha,\gamma}N_{\beta,\delta}+N_{\alpha,\delta}N_{\beta,\gamma})(x_{\alpha}+x_{\beta}+x_{\gamma}-x_{\delta})=0.$
Note that the above computation is not necessary if we use that $dd^{c}\omega=-\partial\overline{\partial}\omega$ is a $(2,2)$-form. On the other hand, for $\alpha,\beta\in\Delta^{+}$, $\gamma,\delta\in\Delta^{-}$, we have that
$dd^{c}\omega(E_{\alpha},E_{\beta},E_{\gamma},E_{\delta})$
$=$ $-N_{\alpha,\beta}N_{\gamma,\delta}(x_{\alpha+\beta}-x_{\gamma}-x_{\delta})+N_{\gamma,\delta}N_{\alpha,\beta}(-x_{\gamma+\delta}+x_{\alpha}+x_{\beta})$
$-\epsilon_{\alpha+\gamma}N_{\alpha,\gamma}N_{\beta,\delta}(\epsilon_{\alpha+\gamma}x_{\alpha+\gamma}+x_{\beta}-x_{\delta})-\epsilon_{\beta+\delta}N_{\beta,\delta}N_{\alpha,\gamma}(\epsilon_{\beta+\delta}x_{\beta+\delta}+x_{\alpha}-x_{\gamma})$
$+\epsilon_{\alpha+\delta}N_{\alpha,\delta}N_{\beta,\gamma}(\epsilon_{\alpha+\delta}x_{\alpha+\delta}+x_{\beta}-x_{\gamma})+\epsilon_{\beta+\gamma}N_{\beta,\gamma}N_{\alpha,\delta}(\epsilon_{\beta+\gamma}x_{\beta+\gamma}+x_{\alpha}-x_{\delta}),$
$=$ $N_{\alpha,\beta}N_{\gamma,\delta}(x_{\alpha}+x_{\beta}+x_{\gamma}+x_{\delta}-2x_{\alpha+\beta})$
$-\epsilon_{\alpha+\gamma}N_{\alpha,\gamma}N_{\beta,\delta}(-x_{\alpha}+x_{\beta}+x_{\gamma}-x_{\delta}+2\epsilon_{\alpha+\gamma}x_{\alpha+\gamma})$
$+\epsilon_{\alpha+\delta}N_{\alpha,\delta}N_{\beta,\gamma}(-x_{\alpha}+x_{\beta}-x_{\gamma}+x_{\delta}+2\epsilon_{\alpha+\delta}x_{\alpha+\delta}).$
Finally, if $\alpha,\beta\in\Delta^{+}$, $\alpha\pm\beta\neq 0$ then
$dd^{c}\omega(E_{\alpha},E_{-\alpha},E_{\beta},E_{-\beta})$
$=$ $-d^{c}\omega(H_{\alpha},E_{\beta},E_{-\beta})+N_{\alpha,\beta}d^{c}\omega(E_{\alpha+\beta},E_{-\alpha},E_{-\beta})-N_{\alpha,-\beta}d^{c}\omega(E_{\alpha-\beta},E_{-\alpha},E_{\beta})$
$-N_{-\alpha,\beta}d^{c}\omega(E_{-\alpha+\beta},E_{\alpha},E_{-\beta})+N_{-\alpha,-\beta}d^{c}\omega(E_{-\alpha-\beta},E_{\alpha},E_{\beta})-d^{c}\omega(H_{\beta},E_{\alpha},E_{-\alpha})$
$=$ $-g_{\mathfrak{t}}(H_{\alpha},H_{\beta})+N_{\alpha,\beta}\mathrm{i}\epsilon_{\alpha+\beta}\epsilon_{-\alpha}\epsilon_{-\beta}N_{\alpha+\beta,-\alpha}(y_{\alpha+\beta}+y_{-\alpha}+y_{-\beta})$
$-N_{\alpha,-\beta}\mathrm{i}\epsilon_{\alpha-\beta}\epsilon_{-\alpha}\epsilon_{\beta}N_{\alpha-\beta,-\alpha}(y_{\alpha-\beta}+y_{-\alpha}+y_{\beta})$
$-N_{-\alpha,\beta}\mathrm{i}\epsilon_{-\alpha+\beta}\epsilon_{\alpha}\epsilon_{-\beta}N_{-\alpha+\beta,\alpha}(y_{-\alpha+\beta}+y_{\alpha}+y_{-\beta})$
$+N_{-\alpha,-\beta}\mathrm{i}\epsilon_{-\alpha-\beta}\epsilon_{\alpha}\epsilon_{\beta}N_{-\alpha-\beta,\alpha}(y_{-\alpha-\beta}+y_{\alpha}+y_{\beta})-g_{\mathfrak{t}}(H_{\beta},H_{\alpha}),$
$=$ $-2g_{\mathfrak{t}}(H_{\alpha},H_{\beta})+2N_{\alpha,\beta}\mathrm{i}N_{\alpha+\beta,-\alpha}(y_{\alpha+\beta}+y_{-\alpha}+y_{-\beta})$
$+2N_{\alpha,-\beta}\mathrm{i}\epsilon_{\alpha-\beta}N_{\alpha-\beta,-\alpha}(y_{\alpha-\beta}+y_{-\alpha}+y_{\beta})$
$=$ $-2g_{\mathfrak{t}}(H_{\alpha},H_{\beta})-2N_{\alpha,\beta}^{2}(x_{\alpha+\beta}-x_{\alpha}-x_{\beta})-2N_{\alpha,-\beta}^{2}\epsilon_{\alpha-\beta}(x_{\alpha-\beta}-x_{\alpha}+x_{\beta}),$
concluding the proof. ∎
Recall that the set $\{\mathrm{i}H_{\alpha}:\alpha\in\Delta^{+}\}$ generates $\mathfrak{t}$.
Corollary 3.3.
A left and $\operatorname{Ad}(T)$-invariant Hermitian structure $(J,g)$ on a compact semisimple Lie group $G$ is pluriclosed if and only if for all $\alpha,\beta\in\Delta^{+}$, $\alpha\neq\beta$,
(1) $g_{\mathfrak{t}}(\mathrm{i}H_{\alpha},\mathrm{i}H_{\beta})=N_{\alpha,\beta}^{2}(x_{\alpha+\beta}-x_{\alpha}-x_{\beta})+\epsilon_{\alpha-\beta}N_{\alpha,-\beta}^{2}(\epsilon_{\alpha-\beta}x_{\alpha-\beta}-x_{\alpha}+x_{\beta}),$
and for all $\alpha,\beta\in\Delta^{+}$, $\gamma,\delta\in\Delta^{-}$ such that $\alpha+\beta+\gamma+\delta=0$ and all pairs add nonzero,
$0=$ $N_{\alpha,\beta}N_{\gamma,\delta}(x_{\alpha}+x_{\beta}+x_{\gamma}+x_{\delta}-2x_{\alpha+\beta})$
(2) $-\epsilon_{\alpha+\gamma}N_{\alpha,\gamma}N_{\beta,\delta}(-x_{\alpha}+x_{\beta}+x_{\gamma}-x_{\delta}+2\epsilon_{\alpha+\gamma}x_{\alpha+\gamma})$
$+\epsilon_{\alpha+\delta}N_{\alpha,\delta}N_{\beta,\gamma}(-x_{\alpha}+x_{\beta}-x_{\gamma}+x_{\delta}+2\epsilon_{\alpha+\delta}x_{\alpha+\delta}).$
As a first application of the above characterization of pluriclosed metrics, actually of only condition (1), we provide an alternative proof for [FG1, Theorem 3.1].
Proposition 3.4.
[FG1] For any left and $\operatorname{Ad}(T)$-invariant pluriclosed Hermitian structure $(J,g)$ on a compact semisimple Lie group $G$, one has that $g_{\mathfrak{t}}=g_{b}|_{\mathfrak{t}}$ for some bi-invariant metric $g_{b}$.
Proof.
Recall that $\{\mathrm{i}H_{\alpha}:\alpha\in\Pi\}$ is a basis of $\mathfrak{t}$, where $\Pi\subset\Delta^{+}$ is the set of simple roots. We set
$\langle\alpha,\beta\rangle:=g_{\operatorname{B}}(\mathrm{i}H_{\alpha},\mathrm{i}H_{\beta})=\operatorname{B}(H_{\alpha},H_{\beta}),\qquad g_{\mathfrak{t}}(\alpha,\beta):=g_{\mathfrak{t}}(\mathrm{i}H_{\alpha},\mathrm{i}H_{\beta}),\quad\forall\alpha,\beta\in\Delta.$
For any $\alpha,\beta\in\Pi$ such that $\langle\alpha,\beta\rangle=0$, we obtain from (1) that $g_{\mathfrak{t}}(\alpha,\beta)=0=\langle\alpha,\beta\rangle$ by using that $\alpha\pm\beta\notin\Delta$. In particular, the maximal tori of two different simple factors are $g_{\mathfrak{t}}$-orthogonal, so we can assume that $G$ is simple and show that $g_{\mathfrak{t}}=g_{\operatorname{B}}|_{\mathfrak{t}}$ up to scaling.
On the other hand, if $\alpha,\beta\in\Pi$ satisfy that $\langle\alpha,\beta\rangle<0$, then there are exactly three possibilities (up to scaling we can assume that $g_{\mathfrak{t}}(\alpha,\alpha)=\langle\alpha,\alpha\rangle$):
  1. (a)
[…]
    $g_{\mathfrak{t}}(\alpha,\beta)=$ $\tfrac{1}{2}\langle\alpha,\alpha\rangle(x_{\alpha+\beta}-x_{\alpha}-x_{\beta}),$
    $g_{\mathfrak{t}}(\alpha+\beta,\alpha)=$ $\tfrac{1}{2}\langle\alpha,\alpha\rangle(x_{\beta}-x_{\alpha+\beta}+x_{\alpha}),$
    $g_{\mathfrak{t}}(\alpha+\beta,\beta)=$ $\tfrac{1}{2}\langle\alpha,\alpha\rangle(x_{\alpha}-x_{\alpha+\beta}+x_{\beta}),$
    which implies that $g_{\mathfrak{t}}(\alpha,\beta)=\langle\alpha,\beta\rangle$, $g_{\mathfrak{t}}(\beta,\beta)=\langle\beta,\beta\rangle$.
  2. (b)
[…]
    $g_{\mathfrak{t}}(\alpha,\beta)=$ $\langle\alpha,\alpha\rangle(x_{\alpha+\beta}-x_{\alpha}-x_{\beta}),$
    $g_{\mathfrak{t}}(\alpha+\beta,\alpha)=$ $\langle\alpha,\alpha\rangle(x_{2\alpha+\beta}-x_{\alpha+\beta}-x_{\alpha})+\langle\alpha,\alpha\rangle(x_{\beta}-x_{\alpha+\beta}+x_{\alpha}),$
    $g_{\mathfrak{t}}(2\alpha+\beta,\alpha+\beta)=$ $\langle\alpha,\alpha\rangle(x_{\alpha}-x_{2\alpha+\beta}+x_{\alpha+\beta}).$
    It is easy to see that the only solutions are $g_{\mathfrak{t}}(\alpha,\beta)=\langle\alpha,\beta\rangle$, $g_{\mathfrak{t}}(\beta,\beta)=\langle\beta,\beta\rangle$.
  3. (c)
[…]
All this allows us to proceed as follows. We start at any end $\alpha_{1}$ of the Dynkin diagram and assume up to scaling that $g_{\mathfrak{t}}(\alpha_{1},\alpha_{1})=\langle\alpha_{1},\alpha_{1}\rangle$. Then, by using (a), (b) and (c) above, we run through the diagram obtaining in each step that $g_{\mathfrak{t}}|_{\operatorname{span}\{\alpha_{i},\alpha_{i+1}\}}=\langle\cdot,\cdot\rangle|_{\operatorname{span}\{\alpha_{i},\alpha_{i+1}\}}$ and the same for the next two nodes in any bifurcation. Thus $g_{\mathfrak{t}}(\mathrm{i}H_{\alpha},\mathrm{i}H_{\beta})=g_{\operatorname{B}}(\mathrm{i}H_{\alpha},\mathrm{i}H_{\beta})$ for any $\alpha,\beta\in\Pi$, that is, $g_{\mathfrak{t}}=g_{\operatorname{B}}|_{\mathfrak{t}}$, as was to be shown. ∎
We are now ready to prove the main result of this paper, a complete classification of left and $\operatorname{Ad}(T)$-invariant pluriclosed metrics on $G$. If $G=G_{1}\times\dots\times G_{s}$, where $G_{i}$ is simple for all $i$, then for each $1\leq i\leq s$ we consider the decomposition $\mathfrak{g}_{i}=\mathfrak{t}_{i}\oplus\mathfrak{q}_{i}$ such that $\mathfrak{t}=\mathfrak{t}_{1}\oplus\dots\oplus\mathfrak{t}_{s}$ and $\mathfrak{q}=\mathfrak{q}_{1}\oplus\dots\oplus\mathfrak{q}_{s}$, so each $\mathfrak{t}_{i}$ is a maximal abelian subalgebra of $\mathfrak{g}_{i}$ and define a root system $\Delta_{i}$ with $\Delta=\Delta_{1}\sqcup\dots\sqcup\Delta_{s}$.
Theorem 3.5.
On any compact semisimple Lie group $G$, the following conditions on a left and $\operatorname{Ad}(T)$-invariant Hermitian structure $(J,g)$ are equivalent:
  1. (i)
    The Hermitian manifold $(G,J,g)$, $g=g_{\mathfrak{t}}+g_{\mathfrak{q}}$ is pluriclosed.
  2. (ii)
    $g_{\mathfrak{t}}(\mathfrak{t}_{i},\mathfrak{t}_{j})=0$ for all $i\neq j$ and for each $1\leq i\leq s$, there exists $z_{i}>0$ such that $g_{\mathfrak{t}}|_{\mathfrak{t}_{i}}=z_{i}g_{\operatorname{B}}|_{\mathfrak{t}_{i}}$ and $g_{\mathfrak{q}}|_{\mathfrak{q}_{i}}=(z_{i}x_{\alpha})_{\alpha\in\Delta^{+}_{i}}$, where the $x_{\alpha}$ are defined as follows: if $\Pi=\{\alpha_{1},\dots,\alpha_{2d_{i}}\}$ are the simple roots of $\mathfrak{g}_{i}$ and we set $x_{j}:=x_{\alpha_{j}}$ for $j=1,\dots,2d_{i}$, then
    $x_{\alpha}:=\sum_{j=1}^{2d_{i}}k_{j}x_{j}-\left(\sum_{j=1}^{2d_{i}}k_{j}-1\right),\qquad\forall\alpha=\sum_{j=1}^{2d_{i}}k_{j}\alpha_{j}\in\Delta^{+}_{i},\quad k_{j}\in{\mathbb{Z}}_{\geq 0}.$
  3. (iii)
    $g_{\mathfrak{t}}(\mathfrak{t}_{i},\mathfrak{t}_{j})=0$ for all $i\neq j$ and on each simple factor, up to scaling, $g_{\mathfrak{t}}=g_{\operatorname{B}}|_{\mathfrak{t}}$ and $x_{\alpha+\beta}=x_{\alpha}+x_{\beta}-1$ for all $\alpha,\beta\in\Delta^{+}$, where $g_{\mathfrak{q}}=(x_{\alpha})_{\alpha\in\Delta^{+}}$.
Remark 3.6.
Given any of these metrics $g$, the space of all left-invariant complex structures which are compatible with $g$ depends on $d(d-1)$ parameters, where $2d=\operatorname{rk}(G)$. On the other hand, it follows from the theorem that for any fixed irreducible $J$, the space of all pluriclosed metrics compatible with $J$ depends on $12d+$ 1 parameters and contains a unique bi-invariant metric up to scaling, given by $g_{b}=z_{1}(-\operatorname{B}_{\mathfrak{g}_{1}})+\dots+z_{s}(-\operatorname{B}_{\mathfrak{g}_{s}})$. Note that $g_{\mathfrak{t}}=g_{b}|_{\mathfrak{t}}$ for any metric $g$ as in the theorem. In general, the number of parameters is $r(2d+1)$ and $J$ admits an $r$-parametric space of Hermitian bi-invariant metrics, where $r$ is the number of irreducible factors of $(G,J)$ as a complex manifold.
Remark 3.7.
The natural number $h(\alpha):=\sum\limits_{i=1}^{n}k_{i}$ is called the height of the root $\alpha$. In order to obtain a metric it is of course necessary that $x_{\alpha}>0$ for all $\alpha\in\Delta^{+}$, which for instance holds if $x_{i}\geq 1-\tfrac{1}{h(\alpha_{max})}$ for all $i=1,\dots,n$, where $\alpha_{max}$ is the maximal root.
Proof.
We first prove that (i) implies (ii). Note that in the proof of Proposition 3.4, we also obtain assuming that $g_{\mathfrak{t}}=g_{\operatorname{B}}|_{\mathfrak{t}}$ the following: $x_{\alpha+\beta}=x_{\alpha}+x_{\beta}-1$ in cases (a) and (b), and $x_{2\alpha+\beta}=2x_{\alpha}+x_{\beta}-2$ in case (b). In case (c), condition (1) gives the following equations when applied to the pairs $(\alpha,\beta)$, $(\alpha+\beta,\alpha)$, $(2\alpha+\beta,\alpha)$ and $(3\alpha+\beta,\beta)$, respectively:
$-1=$ $x_{\alpha+\beta}-x_{\alpha}-x_{\beta},$
$-1=$ $4x_{2\alpha+\beta}-7x_{\alpha+\beta}+3x_{\beta}-x_{\alpha},$
$1=$ $5x_{3\alpha+\beta}-11x_{2\alpha+\beta}+6x_{\alpha+\beta}+x_{\alpha},$
$-1=$ $x_{3\alpha+2\beta}-x_{3\alpha+\beta}-x_{\beta}.$
This implies that $x_{\alpha+\beta}=x_{\alpha}+x_{\beta}-1$, $x_{2\alpha+\beta}=2x_{\alpha}+x_{\beta}-2$, $x_{3\alpha+\beta}=3x_{\alpha}+x_{\beta}-3$ and $x_{3\alpha+2\beta}=3x_{\alpha}+2x_{\beta}-4$. Summarizing, the formula in part (ii) holds for any two positive roots in $\operatorname{span}\{\alpha,\beta\}$, where $\alpha,\beta\in\Pi$.
We now proceed by induction in the height of the roots. A straightforward inspection of the nine types of Dynkin diagrams gives that for any $\gamma=\sum\limits_{i=1}^{2d}k_{i}\alpha_{i}\in\Delta^{+}$ such that at least three different integers $k_{i}$’s are $>0$, there exists $\alpha,\beta\in\Delta^{+}$ such that $\gamma=\alpha+\beta$ and $\alpha-\beta\notin\Delta$. Thus $N_{\alpha,\beta}^{2}=-\langle\alpha,\beta\rangle$ and so by (1), $x_{\gamma}=x_{\alpha+\beta}=x_{\alpha}+x_{\beta}-1$. If $\alpha=\sum\limits_{i=1}^{2d}l_{i}\alpha_{i}$ and $\beta=\sum\limits_{i=1}^{2d}m_{i}\alpha_{i}$ (so $k_{i}=l_{i}+m_{i}$), then by induction hypothesis,
$x_{\gamma}=\sum_{i=1}^{2d}l_{i}x_{i}-(h(\alpha)-1)+\sum_{i=1}^{2d}m_{i}x_{i}-(h(\beta)-1)-1=\sum_{i=1}^{2d}k_{i}x_{i}-(h(\gamma)-1),$
as was to be shown. It is easy to check that part (iii) follows from part (ii).
Finally, if part (iii) holds, then we also have that $x_{\alpha-\beta}+1=\pm(x_{\alpha}-x_{\beta})$ if $\alpha-\beta\in\pm\Delta^{+}$ for all $\alpha,\beta\in\Delta^{+}$. Let us prove part (i) by using Corollary 3.3. We can assume that $G$ is simple and that $g_{\mathfrak{t}}=g_{\operatorname{B}}|_{\mathfrak{t}}$, i.e., $g_{\mathfrak{t}}(\mathrm{i}H_{\alpha},\mathrm{i}H_{\beta})=\langle\alpha,\beta\rangle$ for all $\alpha,\beta\in\Delta^{+}$. By using that $N_{\alpha,-\beta}^{2}=N_{\alpha,\beta}^{2}+\langle\alpha,\beta\rangle$, we obtain that condition (1) for $\alpha,\beta\in\Delta^{+}$, $\alpha\neq\beta$ is equivalent to
$\langle\alpha,\beta\rangle=$ $N_{\alpha,\beta}^{2}(x_{\alpha+\beta}-x_{\alpha}-x_{\beta}+x_{\alpha-\beta}+\epsilon_{\alpha-\beta}(-x_{\alpha}+x_{\beta}))+\langle\alpha,\beta\rangle(x_{\alpha-\beta}+\epsilon_{\alpha-\beta}(-x_{\alpha}+x_{\beta}))$
$=$ $N_{\alpha,\beta}^{2}(1+x_{\alpha-\beta}-(x_{\alpha-\beta}+1))+\langle\alpha,\beta\rangle(x_{\alpha-\beta}-(x_{\alpha-\beta}+1))=\langle\alpha,\beta\rangle,$
so it holds. Concerning condition (2), we have that
$x_{\alpha}+x_{\beta}+x_{\gamma}+x_{\delta}-2x_{\alpha+\beta}=x_{\alpha+\beta}+1+x_{\gamma+\delta}+1-2x_{\alpha+\beta}=2.$
If $\alpha+\gamma\in\Delta^{+}$, then by using that $\alpha=(\alpha+\gamma)+(-\gamma)$ and $-\delta=\beta+(-(\beta+\delta))$, we obtain
$\epsilon_{\alpha+\gamma}(-x_{\alpha}+x_{\beta}+x_{\gamma}-x_{\delta}+2\epsilon_{\alpha+\gamma}x_{\alpha+\gamma})=-x_{\alpha+\gamma}+1-x_{\delta+\beta}+1+2x_{\alpha+\gamma}=2.$
Analogously, it is equal to $-2$ whenever $\alpha+\gamma\in\Delta^{-}$. On the other hand, if $\alpha+\delta\in\Delta^{+}$, then
$\epsilon_{\alpha+\delta}(-x_{\alpha}+x_{\beta}-x_{\gamma}+x_{\delta}+2\epsilon_{\alpha+\delta}x_{\alpha+\delta})=-x_{\alpha+\delta}+1-x_{\beta+\gamma}+1+2x_{\alpha+\delta}=2,$
and it is equal to $-2$ if $\alpha+\delta\in\Delta^{-}$. This implies that (2) is equivalent to
$N_{\alpha,\beta}N_{\gamma,\delta}-N_{\alpha,\gamma}N_{\beta,\delta}+N_{\alpha,\delta}N_{\beta,\gamma}=0,$
which is well known to hold (see §4). Part (i) therefore holds, concluding the proof. ∎
4. Appendix: Roots
The following basic properties of root systems can be found in, among many other books, [H, Chapter III, §4,5, Chapter X, §3] and [K, Chapter II, §5].
Given a connected and compact semisimple Lie group $G$ and a maximal torus $T\subset G$, we have that the Lie algebra $\mathfrak{t}$ of $T$ is a maximal abelian subalgebra of $\mathfrak{g}$, the Lie algebra of $G$, and the complexification $\mathfrak{t}^{c}$ is a Cartan subalgebra of $\mathfrak{g}^{c}$. The adjoint action of $\mathfrak{t}^{c}$ on the complex semisimple Lie algebra $\mathfrak{g}^{c}$ determines the root space decomposition:
$\mathfrak{g}^{c}=\mathfrak{t}^{c}\oplus\bigoplus_{\alpha\in\Delta}\mathfrak{g}_{\alpha},$
where $\Delta\subset(\mathfrak{t}^{c})^{*}$ is the root system, i.e., $[H,E]=\alpha(H)E$ for all $H\in\mathfrak{t}^{c}$ and $E\in\mathfrak{g}_{\alpha}$. Since the Killing form $\operatorname{B}$ of $\mathfrak{g}^{c}$ is non-degenerate on $\mathfrak{t}^{c}$, for any $\alpha\in\Delta$, there is a unique $H_{\alpha}\in\mathfrak{t}^{c}$ such that $\alpha=\operatorname{B}(H_{\alpha},\cdot)$. There exist vectors $E_{\alpha}$, $\alpha\in\Delta$ such that
$\mathfrak{g}_{\alpha}={\mathbb{C}}E_{\alpha},\qquad\operatorname{B}(E_{\alpha},E_{-\alpha})=1,\qquad H_{\alpha}=[E_{\alpha},E_{-\alpha}],\qquad[E_{\alpha},E_{\beta}]=N_{\alpha,\beta}E_{\alpha+\beta}.$
If $\Delta^{+}$ is an ordering of $\Delta$ (i.e., $\Delta=\Delta^{+}\cup-\Delta^{+}$ (disjoint union) and $(\Delta^{+}+\Delta^{+})\cap\Delta\subset\Delta^{+}$) and $\Pi\subset\Delta^{+}$ is the subset of simple roots (i.e., those which are not the sum of two positive roots), then the set
$\{H_{\alpha}:\alpha\in\Pi\}\cup\{E_{\alpha}:\alpha\in\Delta\}$
is a basis of $\mathfrak{g}^{c}$ called the Chevalley basis and the set $\{\mathrm{i}H_{\alpha}:\alpha\in\Pi\}$ is a basis of $\mathfrak{t}$. Every $\beta\in\Delta^{+}$ can be written in a unique way as $\beta=\sum\limits_{\alpha\in\Pi}n_{\alpha}\alpha$, where $n_{\alpha}\in{\mathbb{Z}}_{\geq 0}$ for any $\alpha\in\Pi$, and there exists a unique maximal root $\alpha_{max}$, i.e., for any $\alpha\in\Delta^{+}$, either $\alpha_{\max}-\alpha\in\Delta^{+}$ or $\alpha_{\max}-\alpha\notin\Delta$.
Since $H_{\alpha}\in\mathrm{i}\mathfrak{t}$ for any $\alpha\in\Delta$ and $\operatorname{B}$ is positive definite on $\mathrm{i}\mathfrak{t}$, we can define
$\langle\alpha,\beta\rangle:=\operatorname{B}(H_{\alpha},H_{\beta})=-\operatorname{B}_{\mathfrak{g}}(\mathrm{i}H_{\alpha},\mathrm{i}H_{\beta}),$
where $\operatorname{B}_{\mathfrak{g}}$ is the Killing form of $\mathfrak{g}$.
The numbers $N_{\alpha,\beta}$ such that $[E_{\alpha},E_{\beta}]=N_{\alpha,\beta}E_{\alpha+\beta}$ satisfy the following properties:
$N_{-\alpha,-\beta}=-N_{\alpha,\beta}=N_{\beta,\alpha},\qquad N_{\alpha,\beta}=N_{\beta,\gamma}=N_{\gamma,\alpha},\qquad\mbox{if}\quad\alpha+\beta+\gamma=0,$
and if $\alpha+\beta+\gamma+\delta=0$ and the sum of any pair is nonzero, then
$N_{\alpha,\beta}N_{\gamma,\delta}-N_{\alpha,\gamma}N_{\beta,\delta}+N_{\alpha,\delta}N_{\beta,\gamma}=0.$
Moreover, an explicit formula for $N_{\alpha,\beta}^{2}$ is given as follows. The only multiples of $\alpha\in\Delta$ which are roots are $0,\pm\alpha$, and given $\alpha,\beta\in\Delta$, we have that $\beta+n\alpha\in\Delta$ if and only if $p\leq n\leq q$, where $p\leq 0\leq q$ and
$p+q=-\tfrac{2\langle\alpha,\beta\rangle}{\langle\alpha,\alpha\rangle}\in{\mathbb{Z}},\qquad N_{\alpha,\beta}^{2}=\tfrac{1}{2}q(1-p)\langle\alpha,\alpha\rangle.$
In particular, $N_{\alpha,-\beta}^{2}=N_{\alpha,\beta}^{2}+\langle\alpha,\beta\rangle$, so $N_{\alpha,\beta}^{2}=-\langle\alpha,\beta\rangle$ if $\alpha+\beta\in\Delta$ and $\alpha-\beta\notin\Delta$ (e.g., whenever $\alpha,\beta\in\Pi$).