### v3

# Pluriclosed metrics on compact semisimple Lie groups

Abstract.
Given a compact semisimple Lie group $G$ and a maximal torus $T\subset G$, we give an explicit description of all left and $\operatorname{Ad}(T)$-invariant pluriclosed Hermitian structures on $G$ in terms of the corresponding root system. They depend on $2d+1$ parameters in the irreducible case, where $\dim{T}=2d$. As applications, we obtain that the only left and $\operatorname{Ad}(T)$-invariant pluriclosed metrics which are also CYT are bi-invariant metrics (i.e., Bismut flat) and study the pluriclosed flow as a neat ODE system.
This research was partially supported by three grants from, respectively, CONICET, Univ. Nac. de Córdoba and Foncyt (Argentina)
1. Introduction
A Hermitian metric $g$ on a complex manifold $(M,J)$ is said to be pluriclosed (or SKT, short for strong Kähler with torsion) if the Kähler form $\omega=g(J\cdot,\cdot)$ satisfies that $dd^{c}\omega=0$ (we refer to the recent surveys [F, FG2] for further information). Being $dd^{c}\omega$ a $4$-form, the condition turns out to be very strong in high dimensions, specially in the homogeneous case (see [AL, AN, FP, FS] for structure and classification results on solvmanifolds). The primal compact examples are bi-invariant metrics on any even dimensional compact semisimple Lie group $G$ endowed with any left-invariant complex structure. As known, these are precisely Bismut flat manifolds with finite fundamental group (see [WYZ]). It was proved in [FGV] that a compact homogeneous complex manifold with finite fundamental group (i.e., a C-space) admits an invariant pluriclosed metric only if it is a Lie group (modulo a flag manifold, i.e., a Kähler homogeneous manifold).
Any left-invariant complex structure $J$ on a compact semisimple Lie group $G$ is determined by the choices of a maximal torus $T\subset G$ and of a set of positive roots $\Delta^{+}$ (see [S, W, Pi]), they are often called Samelson complex structures. More precisely, $J$ leaves invariant the $\operatorname{B}_{\mathfrak{g}}$-orthogonal decomposition $\mathfrak{g}=\mathfrak{t}\oplus\mathfrak{q}$ of the Lie algebra $\mathfrak{g}$ of $G$, where $\operatorname{B}_{\mathfrak{g}}$ is the Killing form of $\mathfrak{g}$, $\mathfrak{t}$ is the Lie algebra of $T$, $J_{\mathfrak{t}}$ is any linear map such that $J_{\mathfrak{t}}^{2}=-I$ and on the complexification $\mathfrak{q}^{c}=\bigoplus\limits_{\alpha\in\Delta}{\mathbb{C}}E_{\alpha}$, the complex linear map determined by $J_{\mathfrak{q}}$ on $\mathfrak{q}^{c}$ is given by $J_{\mathfrak{q}}E_{\alpha}=\pm\mathrm{i}E_{\alpha}$ for all $\alpha\in\pm\Delta^{+}$. The space of all left-invariant complex structures on $G$ therefore depends on $2d^{2}$ parameters up to biholomorphism if $\dim{T}=2d$. Any left and $\operatorname{Ad}(T)$-invariant metric which is compatible with a fixed $J=(J_{\mathfrak{t}},J_{\mathfrak{q}})$ is of the form $g=g_{\mathfrak{t}}+g_{\mathfrak{q}}$, where $g_{\mathfrak{t}}$ is any inner product on $\mathfrak{t}$ compatible with $J_{\mathfrak{t}}$ and $g_{\mathfrak{q}}=(x_{\alpha})_{\alpha\in\Delta^{+}}$ is given by
$g_{\mathfrak{q}}=\sum_{\alpha\in\Delta^{+}}x_{\alpha}(-\operatorname{B}_{\mathfrak{g}})|_{\mathfrak{q}_{\alpha}},\qquad x_{\alpha}>0,\quad\mathfrak{q}_{\alpha}:=({\mathbb{C}}E_{\alpha}\oplus{\mathbb{C}}E_{-\alpha})\cap\mathfrak{g},$
producing a space of Hermitian metrics that depends on $d^{2}+|\Delta^{+}|$ parameters.
The following strong results on left-invariant pluriclosed metrics on $G$ were obtained by Fino and Grantcharov in [FG1]:
  1. (a)
    For any left and $\operatorname{Ad}(T)$-invariant pluriclosed metric $g$, there exists a bi-invariant metric $g_{b}$ on $G$ such that $g_{\mathfrak{t}}=g_{b}|_{\mathfrak{t}}$. The proof is based on cohomological properties of the torus fibration $G\rightarrow G/T$ over the full flag manifold $G/T$.
  2. (b)
    A $5$-parametric space of pluriclosed metrics on $G=\mathrm{SO}(9)$ is given for certain complex structure $J$.
  3. (c)
    The only left and $\operatorname{Ad}(T)$-invariant pluriclosed metrics which are also CYT (Calabi-Yau with torsion, i.e., the Bismut Ricci form $\rho^{B}$ vanishes, see §4) are bi-invariant metrics.
Explicit examples of pluriclosed left-invariant metrics (non bi-invariant) were also exhibited in [Ph] for $G=\mathrm{G}_{2}$.
Remark 1.1.
It is worth recalling that given a pluriclosed metric $\omega$ on a compact manifold, $\omega+\overline{\partial}\psi+\partial\overline{\psi}$ is also a pluriclosed metric for any sufficiently small $(1,0)$-form $\psi$. Moreover, using that the $(1,1)$-Aeppli cohomology is one-dimensional (see [B, Theorem 3.1]), one obtains that this construction gives all left-invariant pluriclosed metrics on a compact simple Lie group.
In this paper, after giving a formula for $dd^{c}\omega$ in terms of roots for any left and $\operatorname{Ad}(T)$-invariant Hermitian structure on $G$ (see Lemma 3.1), we provide an alternative algebraic proof for the result stated in (a) above (see Proposition 3.6) and then we prove the following complete classification (see Theorem 3.7 for a more detailed statement). We fix a left-invariant complex structure $J$ attached to a maximal torus $T$ and consider the decomposition $\mathfrak{t}=\mathfrak{t}_{1}\oplus\dots\oplus\mathfrak{t}_{s}$, where the $\mathfrak{t}_{i}$ are the corresponding maximal tori of the simple factors of $G$.
Theorem 1.2.
On a compact semisimple Lie group $G$ endowed with a Samelson complex structure $J$, any left and $\operatorname{Ad}(T)$-invariant pluriclosed metric is given as follows: $g_{\mathfrak{t}}(\mathfrak{t}_{i},\mathfrak{t}_{j})=0$ for all $i\neq j$ and on each simple factor, up to scaling, $g_{\mathfrak{t}}=-\operatorname{B}_{\mathfrak{g}}|_{\mathfrak{t}}$ and $g_{\mathfrak{q}}=(x_{\alpha})_{\alpha\in\Delta^{+}}$, where
$x_{\alpha}:=\sum_{i=1}^{n}k^{\alpha}_{i}x_{i}-\left(\sum_{i=1}^{n}k^{\alpha}_{i}-1\right),\qquad\forall\alpha=\sum_{i=1}^{n}k^{\alpha}_{i}\alpha_{i}\in\Delta^{+},\quad k^{\alpha}_{i}\in{\mathbb{Z}}_{\geq 0},$
$\Pi=\{\alpha_{1},\dots,\alpha_{n}\}\subset\Delta^{+}$ is the set of simple roots and $x_{i}:=x_{\alpha_{i}}$ for $i=1,\dots,n$ (equivalently, $x_{\alpha+\beta}=x_{\alpha}+x_{\beta}-1$ for all $\alpha,\beta,\alpha+\beta\in\Delta^{+}$).
If $\operatorname{rk}(G)=2d$, then each of these metrics is compatible with a d(d1)d(d-1)-parametric space of left-invariant complex structures on $G$. According to the theorem, for a fixed $J$, if the space of all left and $\operatorname{Ad}(T)$-invariant pluriclosed metrics compatible with $J$ is non-empty, then it depends on $r(2d+1)$ parameters and contains an $r$-parametric space of bi-invariant metrics, where $r$ is the number of irreducible factors of $(G,J)$ as a complex manifold. In particular, any irreducible $J$ admits at most one compatible bi-invariant metric up to scaling. Note that $J$ can be irreducible on a non-simple $G$ (see Example 2.1).
Intriguingly enough, the last line in the theorem should be compared with the Kähler condition for the Hermitian structure $(J_{\mathfrak{q}},g_{\mathfrak{q}})$ on the flag manifold $G/T$, given by $x_{\alpha+\beta}=x_{\alpha}+x_{\beta}$ for all $\alpha,\beta,\alpha+\beta\in\Delta^{+}$, whose solutions are precisely
$x_{\alpha}:=\sum\limits_{i=1}^{n}k^{\alpha}_{i}x_{i},\qquad\forall\alpha=\sum\limits_{i=1}^{n}k^{\alpha}_{i}\alpha_{i}\in\Delta^{+}.$
We do not know whether there may be some hidden geometric reason for this behavior.
As a first application of the explicit description of pluriclosed metrics given in Theorem 1.2, we provide in §4 an alternative proof for the result stated in (c) above using the convexity properties of the functional
(1) $F(g)=F(x_{1},\dots,x_{n}):=\sum_{\alpha\in\Delta^{+}}x_{\alpha}-\log{x_{\alpha}}.$
Secondly, we apply Theorem 1.2 to study the pluriclosed flow $\tfrac{\partial}{\partial t}\omega=-(\rho^{B})^{1,1}$ for a one-parameter family $\omega(t)=g(t)(J\cdot,\cdot)$ as an ODE system for
$X(t):=(x_{1}(t),\dots,x_{n}(t))\in{\mathbb{R}}^{n}.$
The pluriclosed flow was introduced in [ST] and in the homogeneous case, it has been studied mostly on solvmanifolds (see [AL, FV] and the references therein). On a compact semisimple Lie group $G$, using a general convergence result from [GJS], Barbaro proved in [B] the convergence to a Bismut flat metric starting at any pluriclosed metric (not necessarily left-invariant) by computing the $(1,1)$-Aeppli cohomology of $G$, obtaining as an application the non-existence of pluriclosed and CYT metrics on $G$ other than the bi-invariant ones.
We show in §5 that the metrics $g(t)$ attached to $X(t)$ pluriclosed flow according to the following ODE system:
(2) $x_{j}^{\prime}=\sum_{\alpha\in\Delta^{+},1\leq i\leq n}\left(1-\tfrac{1}{x_{\alpha}}\right)k^{\alpha}_{i}\langle\alpha_{i},\alpha_{j}\rangle,\qquad\forall j=1,\dots,n,$
or equivalently,
(3) $X^{\prime}(t)=-Q\operatorname{grad}(F)_{X(t)},\qquad\mbox{where}\quad Q:=[\langle\alpha_{i},\alpha_{j}\rangle],$
and $F$ is the functional defined in (1). Thus $Y(t):=Q^{-\tfrac{1}{2}}X(t)$ is the negative gradient flow of the real analytic function $H(Y):=F(Q^{\tfrac{1}{2}}Y)$, which implies that any solution $X(t)$ converges to $(1,\dots,1)$ as $t\to\infty$. Any left and $\operatorname{Ad}(T)$-invariant pluriclosed solution $g(t)$ therefore converges, as $t\to\infty$, to a bi-invariant metric (see Theorem 5.1 for a more precise statement). We note that the ODE systems (2) and (3) clearly describe the pluriclosed flow evolution among a large class of metrics. This may for instance be used to establish the evolution behavior of many geometric quantities along these pluriclosed flow solutions.
Acknowledgements. We thank Beatrice Brienza and Ramiro Lafuente for fruitful conversations and to Fabio Podestà for bringing the paper [AD] to our attention. We are specially grateful to Joseph Kwong for pointing out a mistake in the proof of Theorem 3.7 in a previous version of this paper, and for helping us to fix it. We also thank the anonymous referee for Remark 1.1.
2. Hermitian geometry of Lie groups
Let $G$ be a connected and compact semisimple Lie group of even dimension. The rank of $G$ is therefore also even, say $\operatorname{rk}(G)=2d$. We refer to §6 for the definitions and notation concerning the root system of $G$.
It was proved by Pittie [Pi] that any left-invariant complex structure on $G$ is one of the found independently by Samelson [S] and Wang [W], which are defined as follows. Given a maximal torus $T\subset G$ and an ordering $\Delta^{+}$ of roots, we consider $J:\mathfrak{g}\rightarrow\mathfrak{g}$ given by: the $\operatorname{B}_{\mathfrak{g}}$-orthogonal decomposition $\mathfrak{g}=\mathfrak{t}\oplus\mathfrak{q}$ is $J$-invariant, $J_{\mathfrak{t}}:=J|_{\mathfrak{t}}$ is any linear map such that $J_{\mathfrak{t}}^{2}=-I$ and if
$\mathfrak{g}^{c}=\mathfrak{t}^{c}\oplus\mathfrak{q}^{c},\qquad\mathfrak{q}^{c}=\bigoplus_{\alpha\in\Delta}{\mathbb{C}}E_{\alpha},$
then the complex linear map determined by $J_{\mathfrak{q}}:=J|_{\mathfrak{q}}$ on $\mathfrak{q}^{c}$ is given by
$J_{\mathfrak{q}}E_{\alpha}=\mathrm{i}E_{\alpha},\quad\forall\alpha\in\Delta^{+},\qquad J_{\mathfrak{q}}E_{\alpha}=-\mathrm{i}E_{\alpha},\quad\forall\alpha\in\Delta^{-}:=-\Delta^{+}.$
The biholomorphism class of the complex manifold $(G,J)$, $J=(J_{\mathfrak{t}},J_{\mathfrak{q}})$ does not depend on the choice of $T$ nor of $\Delta^{+}$, so we can fix $J_{\mathfrak{q}}$. Two different $J_{\mathfrak{t}}$’s produce biholomorphic complex structures if and only if they belong to the same orbit of certain finite subgroup of $\operatorname{Aut}(G)$ (see [Pi]). This implies that the space $\mathcal{C}^{G}$ of all left-invariant complex structures on $G$ depends on $2d^{2}$ parameters up to biholomorphism (recall that $\dim{T}=2d$).
The left and $\operatorname{Ad}(T)$-invariant metrics which are compatible with $J=(J_{\mathfrak{t}},J_{\mathfrak{q}})$ are all of the form $g=g_{\mathfrak{t}}+g_{\mathfrak{q}}$, where $g_{\mathfrak{t}}$ is any inner product on $\mathfrak{t}$ compatible with $J_{\mathfrak{t}}$ and $g_{\mathfrak{q}}=(x_{\alpha})_{\alpha\in\Delta}$ is given by
$g_{\mathfrak{q}}=\sum_{\alpha\in\Delta^{+}}x_{\alpha}(-\operatorname{B}_{\mathfrak{g}})|_{\mathfrak{q}_{\alpha}},\qquad\mathfrak{q}_{\alpha}:=({\mathbb{C}}E_{\alpha}\oplus{\mathbb{C}}E_{-\alpha})\cap\mathfrak{g}.$
Equivalently,
$\left\{\tfrac{1}{\sqrt{x_{\alpha}}}e_{\alpha},\;\tfrac{1}{\sqrt{x_{\alpha}}}J_{\mathfrak{q}}e_{\alpha}:\alpha\in\Delta^{+}\right\}$
is a $g$-orthonormal basis of $\mathfrak{q}$. In particular, $x_{\alpha}>0$ and we set $x_{-\alpha}:=x_{\alpha}$ for any $\alpha\in\Delta^{+}$. The corresponding symmetric $\mathbb{C}$}-bilinear form on $\mathfrak{g}^{c}$, also denoted by $g$, satisfies that
$g(E_{\alpha},E_{-\alpha})=-x_{\alpha},\qquad\forall\alpha\in\Delta,$
and $g(E_{\alpha},E_{\beta})=0$ whenever $\alpha+\beta\neq 0$. The Killing metric $g_{\operatorname{B}}:=-\operatorname{B}_{\mathfrak{g}}$ has $x_{\alpha}=1$ for any $\alpha\in\Delta^{+}$ and a bi-invariant metric $g_{b}=z_{1}(-\operatorname{B}_{\mathfrak{g}_{1}})+\dots+z_{s}(-\operatorname{B}_{\mathfrak{g}_{s}})$ corresponds to $x_{\alpha}=z_{i}$ if and only if $\alpha\in\Delta^{+}_{i}$, where $G=G_{1}\times\dots\times G_{s}$ is the decomposition in simple factors and $\Delta_{i}$ is the root system of $G_{i}$.
We note that $(G,J)$ is irreducible as a complex manifold if and only if $J_{\mathfrak{t}}$ does not leave any subspace $\mathfrak{t}_{i_{1}}\oplus\dots\oplus\mathfrak{t}_{i_{k}}$ with $0<k<s$ invariant, where $\mathfrak{t}_{i}$ is the maximal torus of $\mathfrak{g}_{i}$ (cf. [B, Section 2]). In that case, there exists a unique bi-invariant metric compatible with $J$ (up to scaling), which is not necessarily the Killing metric as the following example shows.
Example 2.1.
Consider $G=G_{1}\times G_{2}$ such that $G_{1},G_{2}$ are simple, $\operatorname{rk}(G_{1})=\operatorname{rk}(G_{2})$ and relative to the decomposition $\mathfrak{t}=\mathfrak{t}_{1}\oplus\mathfrak{t}_{2}$,
$J_{\mathfrak{t}}=\left[\begin{matrix}0&-\tfrac{1}{b}I\\ bI&0\end{matrix}\right],\qquad b\neq 0.$
Then $J$ is irreducible and a bi-invariant metric $g_{b}=z_{1}(-\operatorname{B}_{\mathfrak{g}_{1}})+z_{2}(-\operatorname{B}_{\mathfrak{g}_{s}})$ is compatible with $J$ if and only if $\tfrac{z_{1}}{z_{2}}=b$.
We consider the spaces $\mathcal{M}^{T}$, $\mathcal{C}^{T}$, $\mathcal{H}^{T}_{J}$, $\mathcal{C}^{T}_{g}$ and $\mathcal{H}^{T}$ of all left and $\operatorname{Ad}(T)$-invariant metrics, complex structures, metrics compatible with a fixed $J\in\mathcal{C}^{T}$, complex structures compatible with a fixed $g\in\mathcal{M}^{T}$ and Hermitian structures, respectively. They are all manifolds and it is easy to check that their dimensions are respectively given by
$\begin{array}[]{c}\dim{\mathcal{M}^{T}}=\tfrac{d(d+1)}{2}+|\Delta^{+}|,\qquad\dim{\mathcal{C}^{T}}=2d^{2},\\ \\ \dim{\mathcal{H}^{T}_{J}}=d^{2}+|\Delta^{+}|,\qquad\dim{\mathcal{C}^{T}_{g}}=d(d-1),\qquad\dim{\mathcal{H}^{T}}=3d^{2}+|\Delta^{+}|.\end{array}$
The Kähler form $\omega=g(J\cdot,\cdot)$ of a Hermitian structure $(J,g)$ on $G$ as above is given by
$\omega=\omega_{\mathfrak{t}}+\omega_{\mathfrak{q}},\qquad\mbox{where}\quad\omega_{\mathfrak{t}}=g_{\mathfrak{t}}(J_{\mathfrak{t}}\cdot,\cdot),\quad\omega_{\mathfrak{q}}(E_{\alpha},E_{-\alpha})=-\mathrm{i}\epsilon_{\alpha}x_{\alpha},\quad\forall\alpha\in\Delta,$
where $\epsilon_{\alpha}:=\pm 1$ for $\alpha\in\Delta^{\pm}$. Recall that if we consider $H_{\alpha}:=[E_{\alpha},E_{-\alpha}]\in\mathfrak{t}^{c}$, then the set $\{H_{\alpha}:\alpha\in\Delta^{+}\}$ generates $\mathfrak{t}^{c}$.
Lemma 2.2.
For any left and $\operatorname{Ad}(T)$-invariant Hermitian structure $(J,g)$ on a compact semisimple Lie group $G$, the only possibly nonzero components of the $3$-forms $d\omega$ and $d^{c}\omega:=-d\omega(J\cdot,J\cdot,J\cdot)$ are given by
$d\omega(E_{\alpha},E_{\beta},E_{\gamma})=N_{\alpha,\beta}(y_{\alpha}+y_{\beta}+y_{\gamma}),\qquad d^{c}\omega(E_{\alpha},E_{\beta},E_{\gamma})=\mathrm{i}\epsilon_{\alpha}\epsilon_{\beta}\epsilon_{\gamma}N_{\alpha,\beta}(y_{\alpha}+y_{\beta}+y_{\gamma}),$
whenever $\alpha+\beta+\gamma=0$, where $y_{\alpha}=-\mathrm{i}\epsilon_{\alpha}x_{\alpha}$ and $E_{\alpha},E_{\beta}]=N_{\alpha,\beta}E_{\alpha+\beta}$ (see §6), and
$d\omega(A,E_{\alpha},E_{-\alpha})=\omega_{\mathfrak{t}}(A,H_{\alpha})=g_{\mathfrak{t}}(J_{\mathfrak{t}}A,H_{\alpha}),\qquad d^{c}\omega(A,E_{\alpha},E_{-\alpha})=g_{\mathfrak{t}}(A,H_{\alpha}),$
for all $A\in\mathfrak{t}^{c}$.
Remark 2.3.
In particular, $d\omega$ is always nonzero, in accordance with the well-known fact that $(J,g)$ is never Kähler. It also follows that the Hermitian structure $(J_{\mathfrak{q}},g_{\mathfrak{q}})$ on the full flag manifold $G/T$ is Kähler if and only if $x_{\alpha+\beta}=x_{\alpha}+x_{\beta}$ for all $\alpha,\beta,\alpha+\beta\in\Delta^{+}$.
Proof.
For any $\alpha,\beta,\gamma\in\Delta$,
$d\omega(E_{\alpha},E_{\beta},E_{\gamma})$ $=-N_{\alpha,\beta}\omega_{\mathfrak{q}}(E_{\alpha+\beta},E_{\gamma})+N_{\alpha,\gamma}\omega_{\mathfrak{q}}(E_{\alpha+\gamma},E_{\beta})-N_{\beta,\gamma}\omega_{\mathfrak{q}}(E_{\beta+\gamma},E_{\alpha})$
$=N_{\alpha,\beta}y_{\gamma}-N_{\alpha,\gamma}y_{\beta}+N_{\beta,\gamma}y_{\alpha}=N_{\alpha,\beta}(y_{\alpha}+y_{\beta}+y_{\gamma}),$
if $\alpha+\beta+\gamma=0$ and zero otherwise. On the other hand, if $A\in\mathfrak{t}^{c}$, then
$d\omega(A,E_{\alpha},E_{\beta})$ $=-\alpha(A)\omega_{\mathfrak{q}}(E_{\alpha},E_{\beta})+\beta(A)\omega_{\mathfrak{q}}(E_{\beta},E_{\alpha})-\omega([E_{\alpha},E_{\beta}],A)$
$=\left\{\begin{array}[]{l}0,\qquad\alpha+\beta\neq 0,\\ \omega_{\mathfrak{t}}(A,H_{\alpha}),\qquad\alpha+\beta=0.\end{array}\right$.
The only other possibilities, say $d\omega(A,B,E_{\alpha})$ and $d\omega(A,B,C)$ for $B,C\in\mathfrak{t}^{c}$, are easily seen to vanish, concluding the proof. ∎
Left-invariant $1$-forms and closed $2$-forms on $G$ are all respectively given by
$\theta_{X}:=g(\cdot,X),\qquad\sigma_{X}:=\operatorname{B}_{\mathfrak{g}}([\cdot,\cdot],X),\quad X\in\mathfrak{g},$
and using that $\sigma_{PX}=d\theta_{X}$ for all $X$, where $g=g_{\operatorname{B}}(P\cdot,\cdot)$, one obtains the well-known fact that the first and second Betti numbers of $G$ both vanish (see e.g. [LW, Section 2]).
The following vectors in $\mathfrak{t}$ play a crucial role in the geometry of the Hermitian manifold $(G,J,g)$ (see e.g. [AP, G, K, P]):
(4) $Z_{J_{\mathfrak{q}}}:=\sum_{\alpha\in\Delta^{+}}\mathrm{i}H_{\alpha},\qquad Z_{J_{\mathfrak{q}},g_{\mathfrak{q}}}:=\sum_{\alpha\in\Delta^{+}}\tfrac{1}{x_{\alpha}}\mathrm{i}H_{\alpha}.$
Note that $Z_{J_{\mathfrak{q}}}=Z_{J_{\mathfrak{q}},g_{\operatorname{B}}|_{\mathfrak{q}}}$.
Lemma 2.4.
For any left and $\operatorname{Ad}(T)$-invariant Hermitian structure $(J,g)$ on a compact semisimple Lie group $G$,
$d_{g}^{*}\omega=-\theta_{Z_{J_{\mathfrak{q}},g_{\mathfrak{q}}}}\quad\mbox{and}\quad dd_{g}^{*}\omega=-\sigma_{P_{\mathfrak{t}}Z_{J_{\mathfrak{q}},g_{\mathfrak{q}}}},\quad\mbox{where}\quad g_{\mathfrak{t}}=g_{\operatorname{B}}(P_{\mathfrak{t}}\cdot,\cdot).$
Remark 2.5.
In particular, $d_{g}^{*}\omega=-g(\cdot,Z_{J_{\mathfrak{q}},g_{\mathfrak{q}}})$ and
$dd_{g}^{*}\omega=-\operatorname{B}_{\mathfrak{g}}([\cdot,\cdot],P_{\mathfrak{t}}Z_{J_{\mathfrak{q}},g_{\mathfrak{q}}})=g([\cdot,\cdot],Z_{J_{\mathfrak{q}},g_{\mathfrak{q}}}).$
Proof.
For any $X\in\mathfrak{g}$, we have that
$d_{g}^{*}\omega(X)$ $=g(d_{g}^{*}\omega,\theta_{X})=g(\omega,d\theta_{X})=-\sum_{\alpha\in\Delta_{\mathfrak{q}}^{+}}\tfrac{1}{x_{\alpha}^{2}}\omega(e_{\alpha},J_{\mathfrak{q}}e_{\alpha})\theta_{X}([e_{\alpha},J_{\mathfrak{q}}e_{\alpha}])$
$=-\sum_{\alpha\in\Delta_{\mathfrak{q}}^{+}}\tfrac{1}{x_{\alpha}}\theta_{X}(\mathrm{i}H_{\alpha})=-\theta_{X}(Z_{J_{\mathfrak{q}},g_{\mathfrak{q}}})=-g(X,Z_{J_{\mathfrak{q}},g_{\mathfrak{q}}}),$
concluding the proof. ∎
3. Pluriclosed metrics
A Hermitian structure $(J,g)$ is called pluriclosed if $dd^{c}\omega=0$ (or equivalently, $\partial\overline{\partial}\omega=0$). We start this section by computing the $4$-form $dd^{c}\omega$ in our setting.
Lemma 3.1.
For any left and $\operatorname{Ad}(T)$-invariant Hermitian structure $(J,g)$ on a compact semisimple Lie group $G$, the only possibly nonzero components of the $4$-form $dd^{c}\omega$ are given by
$dd^{c}\omega(E_{\alpha},E_{-\alpha},E_{\beta},E_{-\beta})=$ $-2g_{\mathfrak{t}}(H_{\alpha},H_{\beta})-2N_{\alpha,\beta}^{2}(x_{\alpha+\beta}-x_{\alpha}-x_{\beta})$
$-2\epsilon_{\alpha-\beta}N_{\alpha,-\beta}^{2}(\epsilon_{\alpha-\beta}x_{\alpha-\beta}-x_{\alpha}+x_{\beta}),\qquad\forall\alpha,\beta\in\Delta^{+},\alpha\neq\beta,$
and if $\alpha+\beta+\gamma+\delta=0$ and all pairs add nonzero, then for all $\alpha,\beta\in\Delta^{+}$, $\gamma,\delta\in\Delta^{-}$,
$dd^{c}\omega(E_{\alpha},E_{\beta},E_{\gamma},E_{\delta})=$ $N_{\alpha,\beta}N_{\gamma,\delta}(x_{\alpha}+x_{\beta}+x_{\gamma}+x_{\delta}-2x_{\alpha+\beta})$
$-\epsilon_{\alpha+\gamma}N_{\alpha,\gamma}N_{\beta,\delta}(-x_{\alpha}+x_{\beta}+x_{\gamma}-x_{\delta}+2\epsilon_{\alpha+\gamma}x_{\alpha+\gamma})$
$+\epsilon_{\alpha+\delta}N_{\alpha,\delta}N_{\beta,\gamma}(-x_{\alpha}+x_{\beta}-x_{\gamma}+x_{\delta}+2\epsilon_{\alpha+\delta}x_{\alpha+\delta}).$
Remark 3.2.
Using the properties of the numbers $N_{\alpha,\beta}$ given in §6, one can easily check that $g_{\operatorname{B}}$ is therefore always pluriclosed. However, recall that there is a large class of left-invariant complex structures on $G$ which are not compatible with any bi-invariant metric.
Remark 3.3.
These formulas are in accordance with those given in [AD] for full flag manifolds.
Proof.
We use the formula for $d^{c}\omega$ given in Lemma 2.2. If $\alpha+\beta+\gamma=0$, then for any $A\in\mathfrak{t}^{c}$,
$dd^{c}\omega(A,E_{\alpha},E_{\beta},E_{\gamma})$
$=$ $-\alpha(A)d^{c}\omega(E_{\alpha},E_{\beta},E_{\gamma})+\beta(A)d^{c}\omega(E_{\beta},E_{\alpha},E_{\gamma})-\gamma(A)d^{c}\omega(E_{\gamma},E_{\alpha},E_{\beta})$
$-N_{\alpha,\beta}d^{c}\omega(E_{\alpha+\beta},A,E_{\gamma})+N_{\alpha,\gamma}d^{c}\omega(E_{\alpha+\gamma},A,E_{\beta})-N_{\beta,\gamma}d^{c}\omega(E_{\beta+\gamma},A,E_{\alpha})$
$=$ $-(\alpha+\beta+\gamma)(A)d^{c}\omega(E_{\alpha},E_{\beta},E_{\gamma})$
$-N_{\alpha,\beta}d^{c}\omega(A,E_{\gamma},E_{\alpha+\beta})+N_{\alpha,\gamma}d^{c}\omega(A,E_{\beta},E_{\alpha+\gamma})-N_{\beta,\gamma}d^{c}\omega(A,E_{\alpha},E_{\beta+\gamma})$
$=$ $-N_{\alpha,\beta}g_{\mathfrak{t}}(A,H_{\gamma})+N_{\alpha,\gamma}g_{\mathfrak{t}}(A,H_{\beta})-N_{\beta,\gamma}g_{\mathfrak{t}}(A,H_{\alpha})$
$=$ $-N_{\alpha,\beta}g_{\mathfrak{t}}(A,H_{\alpha}+H_{\beta}+H_{\gamma})=-N_{\alpha,\beta}g_{\mathfrak{t}}(A,H_{\alpha+\beta+\gamma})=0.$
For $A,B\in\mathfrak{t}^{c}$ and $\alpha+\beta=0$,
$dd^{c}\omega(A,B,E_{\alpha},E_{\beta})=$ $\alpha(A)d^{c}\omega(E_{\alpha},B,E_{\beta})-\beta(A)d^{c}\omega(E_{\beta},B,E_{\alpha})-\alpha(B)d^{c}\omega(E_{\alpha},A,E_{\beta})$
$+\beta(B)d^{c}\omega(E_{\beta},A,E_{\alpha})-N_{\alpha,\beta}d^{c}\omega(E_{\alpha+\beta},A,B)$
$=$ $-(\alpha+\beta)(A)d^{c}\omega(B,E_{\alpha},E_{\beta})+(\alpha+\beta)(A)d^{c}\omega(A,E_{\alpha},E_{\beta})=0.$
If $\alpha+\beta+\gamma+\delta=0$ and all pairs add nonzero, then
$dd^{c}\omega(E_{\alpha},E_{\beta},E_{\gamma},E_{\delta})$
$=$ $-N_{\alpha,\beta}d^{c}\omega(E_{\alpha+\beta},E_{\gamma},E_{\delta})-N_{\gamma,\delta}d^{c}\omega(E_{\gamma+\delta},E_{\alpha},E_{\beta})$
$+N_{\alpha,\gamma}d^{c}\omega(E_{\alpha+\gamma},E_{\beta},E_{\delta})+N_{\beta,\delta}d^{c}\omega(E_{\beta+\delta},E_{\alpha},E_{\gamma})$
$-N_{\alpha,\delta}d^{c}\omega(E_{\alpha+\delta},E_{\beta},E_{\gamma})-N_{\beta,\gamma}d^{c}\omega(E_{\beta+\gamma},E_{\alpha},E_{\delta})$
$=$ $-N_{\alpha,\beta}\mathrm{i}\epsilon_{\alpha+\beta}\epsilon_{\gamma}\epsilon_{\delta}N_{\gamma,\delta}(y_{\alpha+\beta}+y_{\gamma}+y_{\delta})-N_{\gamma,\delta}\mathrm{i}\epsilon_{\gamma+\delta}\epsilon_{\alpha}\epsilon_{\beta}N_{\alpha,\beta}(y_{\gamma+\delta}+y_{\alpha}+y_{\beta})$
$+N_{\alpha,\gamma}\mathrm{i}\epsilon_{\alpha+\gamma}\epsilon_{\beta}\epsilon_{\delta}N_{\beta,\delta}(y_{\alpha+\gamma}+y_{\beta}+y_{\delta})+N_{\beta,\delta}\mathrm{i}\epsilon_{\beta+\delta}\epsilon_{\alpha}\epsilon_{\gamma}N_{\alpha,\gamma}(y_{\beta+\delta}+y_{\alpha}+y_{\gamma})$
$-N_{\alpha,\delta}\mathrm{i}\epsilon_{\alpha+\delta}\epsilon_{\beta}\epsilon_{\gamma}N_{\beta,\gamma}(y_{\alpha+\delta}+y_{\beta}+y_{\gamma})-N_{\beta,\gamma}\mathrm{i}\epsilon_{\beta+\gamma}\epsilon_{\alpha}\epsilon_{\delta}N_{\alpha,\delta}(y_{\beta+\gamma}+y_{\alpha}+y_{\delta}).$
In particular, for $\alpha,\beta,\gamma\in\Delta^{+}$, $\delta\in\Delta^{-}$, we obtain
$dd^{c}\omega(E_{\alpha},E_{\beta},E_{\gamma},E_{\delta})$
$=$ $N_{\alpha,\beta}N_{\gamma,\delta}(x_{\alpha+\beta}+x_{\gamma}-x_{\delta})+N_{\gamma,\delta}N_{\alpha,\beta}(-x_{\gamma+\delta}+x_{\alpha}+x_{\beta})$
$-N_{\alpha,\gamma}N_{\beta,\delta}(x_{\alpha+\gamma}+x_{\beta}-x_{\delta})-N_{\beta,\delta}N_{\alpha,\gamma}(-x_{\beta+\delta}+x_{\alpha}+x_{\gamma})$
$+N_{\alpha,\delta}N_{\beta,\gamma}(-x_{\alpha+\delta}+x_{\beta}+x_{\gamma})+N_{\beta,\gamma}N_{\alpha,\delta}(x_{\beta+\gamma}+x_{\alpha}-x_{\delta})$
$=$ $(N_{\alpha,\beta}N_{\gamma,\delta}-N_{\alpha,\gamma}N_{\beta,\delta}+N_{\alpha,\delta}N_{\beta,\gamma})(x_{\alpha}+x_{\beta}+x_{\gamma}-x_{\delta})=0.$
Note that the above computation can be skipped by using that $dd^{c}\omega=-\partial\overline{\partial}\omega$ is a $(2,2)$-form. On the other hand, for $\alpha,\beta\in\Delta^{+}$, $\gamma,\delta\in\Delta^{-}$, we have that
$dd^{c}\omega(E_{\alpha},E_{\beta},E_{\gamma},E_{\delta})$
$=$ $-N_{\alpha,\beta}N_{\gamma,\delta}(x_{\alpha+\beta}-x_{\gamma}-x_{\delta})+N_{\gamma,\delta}N_{\alpha,\beta}(-x_{\gamma+\delta}+x_{\alpha}+x_{\beta})$
$-\epsilon_{\alpha+\gamma}N_{\alpha,\gamma}N_{\beta,\delta}(\epsilon_{\alpha+\gamma}x_{\alpha+\gamma}+x_{\beta}-x_{\delta})-\epsilon_{\beta+\delta}N_{\beta,\delta}N_{\alpha,\gamma}(\epsilon_{\beta+\delta}x_{\beta+\delta}+x_{\alpha}-x_{\gamma})$
$+\epsilon_{\alpha+\delta}N_{\alpha,\delta}N_{\beta,\gamma}(\epsilon_{\alpha+\delta}x_{\alpha+\delta}+x_{\beta}-x_{\gamma})+\epsilon_{\beta+\gamma}N_{\beta,\gamma}N_{\alpha,\delta}(\epsilon_{\beta+\gamma}x_{\beta+\gamma}+x_{\alpha}-x_{\delta}),$
$=$ $N_{\alpha,\beta}N_{\gamma,\delta}(x_{\alpha}+x_{\beta}+x_{\gamma}+x_{\delta}-2x_{\alpha+\beta})$
$-\epsilon_{\alpha+\gamma}N_{\alpha,\gamma}N_{\beta,\delta}(-x_{\alpha}+x_{\beta}+x_{\gamma}-x_{\delta}+2\epsilon_{\alpha+\gamma}x_{\alpha+\gamma})$
$+\epsilon_{\alpha+\delta}N_{\alpha,\delta}N_{\beta,\gamma}(-x_{\alpha}+x_{\beta}-x_{\gamma}+x_{\delta}+2\epsilon_{\alpha+\delta}x_{\alpha+\delta}).$
Finally, if $\alpha,\beta\in\Delta^{+}$, $\alpha\pm\beta\neq 0$ then
$dd^{c}\omega(E_{\alpha},E_{-\alpha},E_{\beta},E_{-\beta})$
$=$ $-d^{c}\omega(H_{\alpha},E_{\beta},E_{-\beta})+N_{\alpha,\beta}d^{c}\omega(E_{\alpha+\beta},E_{-\alpha},E_{-\beta})-N_{\alpha,-\beta}d^{c}\omega(E_{\alpha-\beta},E_{-\alpha},E_{\beta})$
$-N_{-\alpha,\beta}d^{c}\omega(E_{-\alpha+\beta},E_{\alpha},E_{-\beta})+N_{-\alpha,-\beta}d^{c}\omega(E_{-\alpha-\beta},E_{\alpha},E_{\beta})-d^{c}\omega(H_{\beta},E_{\alpha},E_{-\alpha})$
$=$ $-g_{\mathfrak{t}}(H_{\alpha},H_{\beta})+N_{\alpha,\beta}\mathrm{i}\epsilon_{\alpha+\beta}\epsilon_{-\alpha}\epsilon_{-\beta}N_{\alpha+\beta,-\alpha}(y_{\alpha+\beta}+y_{-\alpha}+y_{-\beta})$
$-N_{\alpha,-\beta}\mathrm{i}\epsilon_{\alpha-\beta}\epsilon_{-\alpha}\epsilon_{\beta}N_{\alpha-\beta,-\alpha}(y_{\alpha-\beta}+y_{-\alpha}+y_{\beta})$
$-N_{-\alpha,\beta}\mathrm{i}\epsilon_{-\alpha+\beta}\epsilon_{\alpha}\epsilon_{-\beta}N_{-\alpha+\beta,\alpha}(y_{-\alpha+\beta}+y_{\alpha}+y_{-\beta})$
$+N_{-\alpha,-\beta}\mathrm{i}\epsilon_{-\alpha-\beta}\epsilon_{\alpha}\epsilon_{\beta}N_{-\alpha-\beta,\alpha}(y_{-\alpha-\beta}+y_{\alpha}+y_{\beta})-g_{\mathfrak{t}}(H_{\beta},H_{\alpha}),$
$=$ $-2g_{\mathfrak{t}}(H_{\alpha},H_{\beta})+2N_{\alpha,\beta}\mathrm{i}N_{\alpha+\beta,-\alpha}(y_{\alpha+\beta}+y_{-\alpha}+y_{-\beta})$
$+2N_{\alpha,-\beta}\mathrm{i}\epsilon_{\alpha-\beta}N_{\alpha-\beta,-\alpha}(y_{\alpha-\beta}+y_{-\alpha}+y_{\beta})$
$=$ $-2g_{\mathfrak{t}}(H_{\alpha},H_{\beta})-2N_{\alpha,\beta}^{2}(x_{\alpha+\beta}-x_{\alpha}-x_{\beta})-2N_{\alpha,-\beta}^{2}\epsilon_{\alpha-\beta}(\epsilon_{\alpha-\beta}x_{\alpha-\beta}-x_{\alpha}+x_{\beta}),$
concluding the proof. ∎
Recall that the set $\{\mathrm{i}H_{\alpha}:\alpha\in\Delta^{+}\}$ generates $\mathfrak{t}$.
Corollary 3.4.
A left and $\operatorname{Ad}(T)$-invariant Hermitian structure $(J,g)$ on a compact semisimple Lie group $G$ is pluriclosed if and only if for all $\alpha,\beta\in\Delta^{+}$, $\alpha\neq\beta$,
(5) $g_{\mathfrak{t}}(\mathrm{i}H_{\alpha},\mathrm{i}H_{\beta})=N_{\alpha,\beta}^{2}(x_{\alpha+\beta}-x_{\alpha}-x_{\beta})+\epsilon_{\alpha-\beta}N_{\alpha,-\beta}^{2}(\epsilon_{\alpha-\beta}x_{\alpha-\beta}-x_{\alpha}+x_{\beta}),$
and for all $\alpha,\beta\in\Delta^{+}$, $\gamma,\delta\in\Delta^{-}$ such that $\alpha+\beta+\gamma+\delta=0$ and all pairs add nonzero,
$0=$ $N_{\alpha,\beta}N_{\gamma,\delta}(x_{\alpha}+x_{\beta}+x_{\gamma}+x_{\delta}-2x_{\alpha+\beta})$
(6) $-\epsilon_{\alpha+\gamma}N_{\alpha,\gamma}N_{\beta,\delta}(-x_{\alpha}+x_{\beta}+x_{\gamma}-x_{\delta}+2\epsilon_{\alpha+\gamma}x_{\alpha+\gamma})$
$+\epsilon_{\alpha+\delta}N_{\alpha,\delta}N_{\beta,\gamma}(-x_{\alpha}+x_{\beta}-x_{\gamma}+x_{\delta}+2\epsilon_{\alpha+\delta}x_{\alpha+\delta}).$
Remark 3.5.
This characterization coincides with one obtained in the case of a non-compact simple Lie group of inner type in [GP, (4.2)].
As a first application of the above characterization of pluriclosed metrics, actually of only condition (5), we provide an alternative proof for [FG1, Theorem 3.1].
Proposition 3.6.
[FG1] For any left and $\operatorname{Ad}(T)$-invariant pluriclosed Hermitian structure $(J,g)$ on a compact semisimple Lie group $G$, one has that $g_{\mathfrak{t}}=g_{b}|_{\mathfrak{t}}$ for some bi-invariant metric $g_{b}$.
Proof.
Recall that $\{\mathrm{i}H_{\alpha}:\alpha\in\Pi\}$ is a basis of $\mathfrak{t}$, where $\Pi\subset\Delta^{+}$ is the set of simple roots. We set
$\langle\alpha,\beta\rangle:=g_{\operatorname{B}}(\mathrm{i}H_{\alpha},\mathrm{i}H_{\beta})=\operatorname{B}(H_{\alpha},H_{\beta}),\qquad g_{\mathfrak{t}}(\alpha,\beta):=g_{\mathfrak{t}}(\mathrm{i}H_{\alpha},\mathrm{i}H_{\beta}),\quad\forall\alpha,\beta\in\Delta.$
For any $\alpha,\beta\in\Pi$ such that $\langle\alpha,\beta\rangle=0$, we obtain from (5) that $g_{\mathfrak{t}}(\alpha,\beta)=0=\langle\alpha,\beta\rangle$ by using that $\alpha\pm\beta\notin\Delta$. In particular, the maximal tori of two different simple factors are $g_{\mathfrak{t}}$-orthogonal, so we can assume that $G$ is simple and show that $g_{\mathfrak{t}}=g_{\operatorname{B}}|_{\mathfrak{t}}$ up to scaling.
On the other hand, if $\alpha,\beta\in\Pi$ satisfy that $\langle\alpha,\beta\rangle<0$, then there are exactly three possibilities (up to scaling we can assume that $g_{\mathfrak{t}}(\alpha,\alpha)=\langle\alpha,\alpha\rangle$):
  1. (a)
    $\alpha$$\beta$ : $\langle\alpha,\alpha\rangle=\langle\beta,\beta\rangle=-2\langle\alpha,\beta\rangle$, $N_{\alpha,\beta}^{2}=\tfrac{1}{2}\langle\alpha,\alpha\rangle=N_{\alpha+\beta,\alpha}^{2}=N_{\alpha+\beta,\beta}^{2}$. It follows from (5) that
    $g_{\mathfrak{t}}(\alpha,\beta)=$ $\tfrac{1}{2}\langle\alpha,\alpha\rangle(x_{\alpha+\beta}-x_{\alpha}-x_{\beta}),$
    $g_{\mathfrak{t}}(\alpha+\beta,\alpha)=$ $\tfrac{1}{2}\langle\alpha,\alpha\rangle(x_{\beta}-x_{\alpha+\beta}+x_{\alpha}),$
    $g_{\mathfrak{t}}(\alpha+\beta,\beta)=$ $\tfrac{1}{2}\langle\alpha,\alpha\rangle(x_{\alpha}-x_{\alpha+\beta}+x_{\beta}),$
    which implies that $g_{\mathfrak{t}}(\alpha,\beta)=\langle\alpha,\beta\rangle$, $g_{\mathfrak{t}}(\beta,\beta)=\langle\beta,\beta\rangle$.
  2. (b)
    $\alpha$$\beta$ : $\langle\alpha,\alpha\rangle=\tfrac{1}{2}\langle\beta,\beta\rangle=-\langle\alpha,\beta\rangle$, $N_{\alpha,\beta}^{2}=\langle\alpha,\alpha\rangle=N_{\alpha+\beta,\alpha}^{2}=N_{2\alpha+\beta,\alpha+\beta}^{2}$. By (5) we have that $g_{\mathfrak{t}}(2\alpha+\beta,\beta)=0$, so $g_{\mathfrak{t}}(\alpha,\beta)=-\tfrac{1}{2}g_{\mathfrak{t}}(\beta,\beta)$, and in addition,
    $g_{\mathfrak{t}}(\alpha,\beta)=$ $\langle\alpha,\alpha\rangle(x_{\alpha+\beta}-x_{\alpha}-x_{\beta}),$
    $g_{\mathfrak{t}}(\alpha+\beta,\alpha)=$ $\langle\alpha,\alpha\rangle(x_{2\alpha+\beta}-x_{\alpha+\beta}-x_{\alpha})+\langle\alpha,\alpha\rangle(x_{\beta}-x_{\alpha+\beta}+x_{\alpha}),$
    $g_{\mathfrak{t}}(2\alpha+\beta,\alpha+\beta)=$ $\langle\alpha,\alpha\rangle(x_{\alpha}-x_{2\alpha+\beta}+x_{\alpha+\beta}).$
    It is easy to see that the only solutions are $g_{\mathfrak{t}}(\alpha,\beta)=\langle\alpha,\beta\rangle$, $g_{\mathfrak{t}}(\beta,\beta)=\langle\beta,\beta\rangle$.
  3. (c)
    $\beta$$\alpha$ : $\langle\alpha,\alpha\rangle=\tfrac{1}{3}\langle\beta,\beta\rangle=-\tfrac{2}{3}\langle\alpha,\beta\rangle$, $N_{\alpha,\beta}^{2}=\tfrac{3}{2}\langle\alpha,\alpha\rangle=N_{2\alpha+\beta,\alpha}^{2}=N_{3\alpha+\beta,\beta}^{2}$, $N_{\alpha+\beta,\alpha}^{2}=2\langle\alpha,\alpha\rangle$. Here condition (5) gives that $g_{\mathfrak{t}}(2\alpha+\beta,\beta)=0$ and $g_{\mathfrak{t}}(3\alpha+\beta,\alpha+\beta)=0$, from which clearly follows that $g_{\mathfrak{t}}(\alpha,\beta)=\langle\alpha,\beta\rangle$ and $g_{\mathfrak{t}}(\beta,\beta)=\langle\beta,\beta\rangle$.
All this allows us to proceed as follows. We start at any end $\alpha_{1}$ of the Dynkin diagram and assume up to scaling that $g_{\mathfrak{t}}(\alpha_{1},\alpha_{1})=\langle\alpha_{1},\alpha_{1}\rangle$. Then, by using (a), (b) and (c) above, we run through the diagram obtaining in each step that $g_{\mathfrak{t}}|_{\operatorname{span}\{\alpha_{i},\alpha_{i+1}\}}=\langle\cdot,\cdot\rangle|_{\operatorname{span}\{\alpha_{i},\alpha_{i+1}\}}$ and the same for the next two nodes in any bifurcation. Thus $g_{\mathfrak{t}}(\mathrm{i}H_{\alpha},\mathrm{i}H_{\beta})=g_{\operatorname{B}}(\mathrm{i}H_{\alpha},\mathrm{i}H_{\beta})$ for any $\alpha,\beta\in\Pi$, that is, $g_{\mathfrak{t}}=g_{\operatorname{B}}|_{\mathfrak{t}}$, as was to be shown. ∎
We are now ready to prove the main result of this paper, a complete classification of left and $\operatorname{Ad}(T)$-invariant pluriclosed metrics on $G$. If $G=G_{1}\times\dots\times G_{s}$, where $G_{i}$ is simple for all $i$, then for each $1\leq i\leq s$ we consider the decomposition $\mathfrak{g}_{i}=\mathfrak{t}_{i}\oplus\mathfrak{q}_{i}$ such that $\mathfrak{t}=\mathfrak{t}_{1}\oplus\dots\oplus\mathfrak{t}_{s}$ and $\mathfrak{q}=\mathfrak{q}_{1}\oplus\dots\oplus\mathfrak{q}_{s}$, so each $\mathfrak{t}_{i}$ is a maximal abelian subalgebra of $\mathfrak{g}_{i}$ and defines a root system $\Delta_{i}$ with $\Delta=\Delta_{1}\sqcup\dots\sqcup\Delta_{s}$.
Theorem 3.7.
On any compact semisimple Lie group $G$, the following conditions on a left and $\operatorname{Ad}(T)$-invariant Hermitian structure $(J,g)$ are equivalent:
  1. (i)
    The Hermitian manifold $(G,J,g)$, $g=g_{\mathfrak{t}}+g_{\mathfrak{q}}$ is pluriclosed.
  2. (ii)
    $g_{\mathfrak{t}}(\mathfrak{t}_{i},\mathfrak{t}_{j})=0$ for all $i\neq j$ and on each simple factor, up to scaling, $g_{\mathfrak{t}}=g_{\operatorname{B}}|_{\mathfrak{t}}$ and $x_{\alpha+\beta}=x_{\alpha}+x_{\beta}-1$ for all $\alpha,\beta,\alpha+\beta\in\Delta^{+}$, where $g_{\mathfrak{q}}=(x_{\alpha})_{\alpha\in\Delta^{+}}$.
  3. (iii)
    $g_{\mathfrak{t}}(\mathfrak{t}_{i},\mathfrak{t}_{j})=0$ for all $i\neq j$ and for each $1\leq i\leq s$, there exists $z_{i}>0$ such that $g_{\mathfrak{t}}|_{\mathfrak{t}_{i}}=z_{i}g_{\operatorname{B}}|_{\mathfrak{t}_{i}}$ and $g_{\mathfrak{q}}|_{\mathfrak{q}_{i}}=(z_{i}x_{\alpha})_{\alpha\in\Delta^{+}_{i}}$, where the $x_{\alpha}$ are defined as follows: if $\Pi_{i}=\{\alpha_{1},\dots,\alpha_{n_{i}}\}$ are the simple roots of $\mathfrak{g}_{i}$ and we set $x_{j}:=x_{\alpha_{j}}$ for $j=1,\dots,n_{i}$, then x_α:=_j=1^n_i k^α_jx_j -(_j=1^n_i k^α_j -1),   α=_j=1^n_i k^α_jα_jΔ^+_i,  k^α_jZ_0.
Remark 3.8.
Given any of these metrics $g$, the space of all left-invariant complex structures which are compatible with $g$ depends on d(d1)d(d-1) parameters, where $2d=\operatorname{rk}(G)$. On the other hand, it follows from the theorem that for any fixed irreducible $J$, the space of all pluriclosed metrics compatible with $J$, if non-empty, depends on $2d+1$ parameters and contains a unique bi-invariant metric up to scaling, given by $g_{b}=z_{1}(-\operatorname{B}_{\mathfrak{g}_{1}})+\dots+z_{s}(-\operatorname{B}_{\mathfrak{g}_{s}})$. Note that $g_{\mathfrak{t}}=g_{b}|_{\mathfrak{t}}$ for any metric $g$ as in the theorem. In general, the number of parameters is $r(2d+1)$ and $J$ admits an $r$-parametric space of Hermitian bi-invariant metrics, where $r$ is the number of irreducible factors of $(G,J)$ as a complex manifold.
Remark 3.9.
The natural number $h(\alpha):=\sum\limits_{i=1}^{n}k^{\alpha}_{i}$ is called the height of the root $\alpha=\sum\limits_{i=1}^{n}k^{\alpha}_{i}\alpha_{i}$, where $\Pi=\{\alpha_{1},\dots,\alpha_{n}\}$. In order to obtain a metric it is of course necessary that $x_{\alpha}>0$ for all $\alpha\in\Delta^{+}$, which for instance holds if $x_{i}\geq 1-\tfrac{1}{h(\alpha_{max})}$ for all $i=1,\dots,n$, where $\alpha_{max}$ is the maximal root.
Proof.
We first prove that (i) implies (ii) by induction in the height of the roots. On any of the simple factors of $G$, we can assume by Proposition 3.6 that $g_{\mathfrak{t}}(\mathrm{i}H_{\alpha},\mathrm{i}H_{\beta})=\langle\alpha,\beta\rangle$ for all $\alpha,\beta\in\Delta$, so it follows from (5) that
$\langle\alpha,\beta\rangle=N_{\alpha,\beta}^{2}(x_{\alpha+\beta}-x_{\alpha}-x_{\beta})+\epsilon_{\alpha-\beta}N_{\alpha,-\beta}^{2}(\epsilon_{\alpha-\beta}x_{\alpha-\beta}-x_{\alpha}+x_{\beta}),$
for all $\alpha,\beta\in\Delta^{+}$, $\alpha\neq\beta$. If $\epsilon_{\alpha-\beta}=0$, then $N_{\alpha,\beta}^{2}=-\langle\alpha,\beta\rangle$ and we are done. Otherwise, we can assume that $\epsilon_{\alpha-\beta}=1$, i.e., $\alpha-\beta\in\Delta^{+}$, and since $x_{\alpha}=x_{\alpha-\beta}+x_{\beta}-1$ by induction hypothesis, we obtain that (see §6)
$\langle\alpha,\beta\rangle=N_{\alpha,\beta}^{2}(x_{\alpha+\beta}-x_{\alpha}-x_{\beta})+N_{\alpha,\beta}^{2}+\langle\alpha,\beta\rangle,$
so $x_{\alpha+\beta}-x_{\alpha}-x_{\beta}=-1$.
On the other hand, it is easy to check that part (iii) is equivalent to part (ii).
Finally, if part (ii) holds, then we also have that $x_{\alpha-\beta}-1=\pm(x_{\alpha}-x_{\beta})$ if $\alpha-\beta\in\pm\Delta^{+}$ for all $\alpha,\beta\in\Delta^{+}$. Let us prove part (i) by using Corollary 3.4. We can assume that $G$ is simple and that $g_{\mathfrak{t}}=g_{\operatorname{B}}|_{\mathfrak{t}}$, i.e., $g_{\mathfrak{t}}(\mathrm{i}H_{\alpha},\mathrm{i}H_{\beta})=\langle\alpha,\beta\rangle$ for all $\alpha,\beta\in\Delta^{+}$. By using that $N_{\alpha,-\beta}^{2}=N_{\alpha,\beta}^{2}+\langle\alpha,\beta\rangle$, we obtain that condition (5) for $\alpha,\beta\in\Delta^{+}$, $\alpha\neq\beta$ is equivalent to
$\langle\alpha,\beta\rangle=$ $N_{\alpha,\beta}^{2}(x_{\alpha+\beta}-x_{\alpha}-x_{\beta}+x_{\alpha-\beta}+\epsilon_{\alpha-\beta}(-x_{\alpha}+x_{\beta}))+\langle\alpha,\beta\rangle(x_{\alpha-\beta}+\epsilon_{\alpha-\beta}(-x_{\alpha}+x_{\beta}))$
$=$ $N_{\alpha,\beta}^{2}(-1+x_{\alpha-\beta}-(x_{\alpha-\beta}-1))+\langle\alpha,\beta\rangle(x_{\alpha-\beta}-(x_{\alpha-\beta}-1))=\langle\alpha,\beta\rangle,$
so it holds. Concerning condition (6), we have that
$x_{\alpha}+x_{\beta}+x_{\gamma}+x_{\delta}-2x_{\alpha+\beta}=x_{\alpha+\beta}+1+x_{\gamma+\delta}+1-2x_{\alpha+\beta}=2.$
If $\alpha+\gamma\in\Delta^{+}$, then by using that $\alpha=(\alpha+\gamma)+(-\gamma)$ and $\delta=\beta+(-(\beta+\delta))$, we obtain
$\epsilon_{\alpha+\gamma}(-x_{\alpha}+x_{\beta}+x_{\gamma}-x_{\delta}+2\epsilon_{\alpha+\gamma}x_{\alpha+\gamma})=-x_{\alpha+\gamma}+1-x_{\delta+\beta}+1+2x_{\alpha+\gamma}=2.$
Analogously, it is equal to 2-2 whenever $\alpha+\gamma\in\Delta^{-}$. On the other hand, if $\alpha+\delta\in\Delta^{+}$, then
$\epsilon_{\alpha+\delta}(-x_{\alpha}+x_{\beta}-x_{\gamma}+x_{\delta}+2\epsilon_{\alpha+\delta}x_{\alpha+\delta})=-x_{\alpha+\delta}+1-x_{\beta+\gamma}+1+2x_{\alpha+\delta}=2,$
and it is equal to 2-2 if $\alpha+\delta\in\Delta^{-}$. This implies that (6) is equivalent to
$N_{\alpha,\beta}N_{\gamma,\delta}-N_{\alpha,\gamma}N_{\beta,\delta}+N_{\alpha,\delta}N_{\beta,\gamma}=0,$
which is well known to hold (see §6). Part (i) therefore holds, concluding the proof. ∎
Remark 3.10.
The above proof shows that condition (6) actually follows from (5) in the characterization of pluriclosed given in Corollary 3.4.
4. CYT metrics
On any Hermitian manifold $(M^{2n},J,g,\omega)$, there is a line $\nabla^{t}$, $t\in{\mathbb{R}}$ of canonical connections which are Hermitian (i.e., $\nabla^{t}J=0$ and $\nabla^{t}g=0$) and with torsion first studied by Gauduchon in [G] and defined by
$g(\nabla^{t}_{X}Y,Z)=g(\nabla^{g}_{X}Y,Z)+\tfrac{t-1}{4}d^{c}\omega(X,Y,Z)+\tfrac{t+1}{4}d^{c}\omega(X,JY,JZ),$
where $\nabla^{g}$ is the Levi-Civita connection of $(M,g)$. Their corresponding (first) Ricci forms, given by
$\rho^{t}(X,Y):=-\tfrac{1}{2}\sum_{i=1}^{2n}g(R^{t}(X,Y)e_{i},Je_{i}),$
where $\{e_{i}\}_{1}^{2n}$ is a $g$-orthonormal frame and $R^{t}(X,Y):=[\nabla^{t}_{X},\nabla^{t}_{Y}]-\nabla^{t}_{[X,Y]}$ is the curvature of $\nabla^{t}$, are all closed $2$-forms representing the first Chern class $c_{1}(M,J)$ up to scaling.
We are interested in this paper in the Chern connection $\nabla^{C}:=\nabla^{1}$ and the Bismut (or Strominger) connection $\nabla^{B}:=\nabla^{-1}$, whose Ricci forms are related by (see [IP, (2.10)])
(7) $\rho^{B}=\rho^{C}-dd_{g}^{*}\omega.$
The metric $g$ is called Calabi-Yau with torsion (CYT for short) if $\rho^{B}=0$, i.e., the Bismut Ricci form vanishes.
Back in the setting of left-invariant Hermitian structures on Lie groups, according to [V, (4.2)] (see also [LR, (2.2)] and [GP, Section 5]), the Chern Ricci form is given by
$\rho^{C}(X,Y)=-\tfrac{1}{2}\operatorname{tr}{J\operatorname{ad}{[X,Y]}},\qquad\forall X,Y\in\mathfrak{g}.$
This implies that on compact semisimple Lie groups,
$\rho^{C}(X,Y)=$ $-\tfrac{1}{2}\sum_{\alpha\in\Delta^{+}}2\mathrm{i}\alpha([X,Y]_{\mathfrak{t}})=-\sum_{\alpha\in\Delta^{+}}\mathrm{i}\operatorname{B}([X,Y],H_{\alpha})$
$=$ $-\sum_{\alpha\in\Delta^{+}}\operatorname{B}_{\mathfrak{g}}([X,Y],\mathrm{i}H_{\alpha})=-\operatorname{B}_{\mathfrak{g}}([\cdot,\cdot],Z_{J_{\mathfrak{q}}}),$
where $Z_{J_{\mathfrak{q}}}$ is as in (4) (see also [FG1, Section 4] and [AP, K]). We therefore obtain from Lemma 2.4 and (7) that
(8) $\rho^{B}=\operatorname{B}_{\mathfrak{g}}\left([\cdot,\cdot],-Z_{J_{\mathfrak{q}}}+P_{\mathfrak{t}}Z_{J_{\mathfrak{q}},g_{\mathfrak{q}}}\right),\qquad g_{\mathfrak{t}}=g_{\operatorname{B}}(P_{\mathfrak{t}}\cdot,\cdot).$
In particular, the metric $g=g_{\mathfrak{t}}+g_{\mathfrak{q}}$ on the complex manifold $(G,J)$ is CYT if and only if $P_{\mathfrak{t}}Z_{J_{\mathfrak{q}},g_{\mathfrak{q}}}=Z_{J_{\mathfrak{q}}}$. We are now ready to give, as an application of Theorem 3.7, an alternative proof for [FG1, Corollary 5.1, iii)].
Proposition 4.1.
[FG1] A left and $\operatorname{Ad}(T)$-invariant Hermitian structure $(J,g)$ on a compact semisimple Lie group $G$ is pluriclosed and CYT if and only if $g$ is a bi-invariant metric (i.e., Bismut flat).
Proof.
Recall from Theorem 3.7 the form of any left and $\operatorname{Ad}(T)$-invariant pluriclosed metric $g$. We can therefore assume that $G$ is simple and $P_{\mathfrak{t}}=I$, which implies that $g$ is CYT if and only if $Z_{J_{\mathfrak{q}},g_{\mathfrak{q}}}=Z_{J_{\mathfrak{q}}}$ (see (8)), that is,
$\sum_{\alpha\in\Delta^{+}}\left(1-\tfrac{1}{x_{\alpha}}\right)\mathrm{i}H_{\alpha}=0,\quad\mbox{where}\quad x_{\alpha}=1+\sum_{i=1}^{n}k^{\alpha}_{i}(x_{i}-1)\quad\mbox{if}\quad\alpha=\sum_{i=1}^{n}k^{\alpha}_{i}\alpha_{i}.$
Using that $\{\mathrm{i}H_{\alpha_{1}},\dots,\mathrm{i}H_{\alpha_{n}}\}$ is a basis of $\mathfrak{t}$, we obtain that this is equivalent to
(9) $\sum_{\alpha\in\Delta^{+}}\left(1-\tfrac{1}{x_{\alpha}}\right)k^{\alpha}_{j}=0,\qquad\forall j=1,\dots,n.$
We need to show that this implies that $x_{1}=\dots=x_{n}=1$ in order to prove the proposition.
For each fixed nonzero vector $(y_{1},\dots,y_{n})$, we consider the straight line
$x_{i}=x_{i}(t):=1+ty_{i},\qquad i=1,\dots,n,\quad\mbox{so}\quad x_{\alpha}=x_{\alpha}(t)=1+\sum_{i=1}^{n}k^{\alpha}_{i}ty_{i},$
and the function $f(t):=\sum\limits_{\alpha\in\Delta^{+}}x_{\alpha}-\log{x_{\alpha}}$. Thus $f^{\prime}(0)=0$,
$f^{\prime}(t)=\sum\limits_{\alpha\in\Delta^{+}}x_{\alpha}^{\prime}\left(1-\tfrac{1}{x_{\alpha}}\right)=\sum\limits_{i=1}^{n}y_{i}\sum\limits_{\alpha\in\Delta^{+}}k^{\alpha}_{i}\left(1-\tfrac{1}{x_{\alpha}}\right),\qquad f^{\prime\prime}(t)=\sum\limits_{\alpha\in\Delta^{+}}\tfrac{(x_{\alpha}^{\prime})^{2}}{x_{\alpha}^{2}}.$
Since $f^{\prime\prime}(t)>0$ for all $t$ (note that $x_{\alpha_{i}}^{\prime}(t)=y_{i}$ for all $i=1,\dots,n$), we have that $f^{\prime}(t)=0$ if and only if $t=0$. But if the metric $g(x_{1},\dots,x_{n})$ is CYT then $f^{\prime}(t)=0$ by (9), concluding the proof. ∎
5. Pluriclosed flow
On a given complex manifold $(M,J)$, the pluriclosed flow for a one-parameter family of Hermitian metrics $\omega(t)=g(t)(J\cdot,\cdot)$ is defined by
(10) $\tfrac{\partial}{\partial t}\omega=-(\rho^{B})^{1,1},\qquad\omega(0)=\omega_{0},$
where $(\rho^{B})^{1,1}$ is the $(1,1)$-component of the Bismut Ricci form $\rho^{B}$ (see (7)).
According to Theorem 3.7, any left and $\operatorname{Ad}(T)$-invariant pluriclosed metric $g$ on a given compact semisimple Lie group $G$ endowed with a Samelson complex structure $J$ is of the form $g=g_{\mathfrak{t}}+g_{\mathfrak{q}}$, and up to scaling, on each simple factor of $G$ we have that $g_{\mathfrak{t}}=g_{\operatorname{B}}|_{\mathfrak{t}}$, $g_{\mathfrak{q}}=(x_{\alpha})_{\alpha\in\Delta^{+}}$ and if $\Pi=\{\alpha_{1},\dots,\alpha_{n}\}$ and we set $x_{i}:=x_{\alpha_{i}}$, then
(11) $x_{\alpha}=1+\sum_{i=1}^{n}k^{\alpha}_{i}(x_{i}-1),\qquad\forall\alpha=\sum_{i=1}^{n}k^{\alpha}_{i}\alpha_{i}.$
It follows from (8) that the pluriclosed flow evolution for $g(t)=g_{\mathfrak{t}}(t)+g_{\mathfrak{q}}(t)$, $g_{\mathfrak{q}}(t)=(x_{\alpha}(t))_{\alpha\in\Delta^{+}}$ is given on each simple factor of $G$ by $g_{\mathfrak{t}}(t)=g_{\operatorname{B}}|_{\mathfrak{t}}$ for all $t$ and for each $\beta\in\Delta^{+}$,
$-\mathrm{i}x_{\beta}^{\prime}=$ $y_{\beta}^{\prime}=\tfrac{d}{dt}\omega_{\mathfrak{q}}(E_{\beta},E_{-\beta})=-\rho^{B}(E_{\beta},E_{-\beta})=\operatorname{B}_{\mathfrak{g}}(H_{\beta},Z_{J_{\mathfrak{q}}}-Z_{J_{\mathfrak{q}},g_{\mathfrak{q}}})$
$=-\mathrm{i}\operatorname{B}_{\mathfrak{g}}\left(\mathrm{i}H_{\beta},\sum_{\alpha\in\Delta^{+}}\left(1-\tfrac{1}{x_{\alpha}}\right)\mathrm{i}H_{\alpha}\right)=\mathrm{i}\sum_{\alpha\in\Delta^{+}}\left(1-\tfrac{1}{x_{\alpha}}\right)g_{\operatorname{B}}(\mathrm{i}H_{\beta},\mathrm{i}H_{\alpha}),$
which is equivalent to
(12) $x_{\beta}^{\prime}=-\sum_{\alpha\in\Delta^{+}}\left(1-\tfrac{1}{x_{\alpha}}\right)\langle\alpha,\beta\rangle,\qquad\forall\beta\in\Delta^{+}.$
By (11), the pluriclosed flow (10) is therefore equivalent to the following ODE system:
(13) $x_{j}^{\prime}=-\sum_{\alpha\in\Delta^{+},1\leq i\leq n}\left(1-\tfrac{1}{x_{\alpha}}\right)k^{\alpha}_{i}\langle\alpha_{i},\alpha_{j}\rangle,\qquad\forall j=1,\dots,n.$
We consider the convex functional already used in §4 given by
$F(X):=\sum_{\alpha\in\Delta^{+}}x_{\alpha}-\log{x_{\alpha}},\qquad\forall X:=(x_{1},\dots,x_{n}),$
which has
$\tfrac{\partial}{\partial x_{i}}F=\sum_{\alpha\in\Delta^{+}}k^{\alpha}_{i}\left(1-\tfrac{1}{x_{\alpha}}\right),\qquad\forall i=1,\dots,n.$
See Remark 3.9 about the positivity region of the $x_{\alpha}$. In the proof of Proposition 4.1, we showed that $F(X)>F(X_{0})$ for all $X\neq X_{0}:=(1,\dots,1)$.
It follows from (9) and Proposition 4.1 that $\operatorname{grad}(F)_{X}=0$ if and only if $X=X_{0}$, where $\operatorname{grad}(F)_{X}=(\tfrac{\partial}{\partial x_{1}}|_{X}F,\dots,\tfrac{\partial}{\partial x_{n}}|_{X}F)$ is the gradient of $F$ at the point $X$. Using (13), we obtain that the pluriclosed flow can be written as
(14) $X^{\prime}(t)=-Q\operatorname{grad}(F)_{X(t)},\qquad\mbox{where}\quad Q:=[\langle\alpha_{i},\alpha_{j}\rangle].$
Note that $Q$ is a positive definite symmetric $n\times n$ matrix. Thus
$\tfrac{d}{dt}F(X(t))=\langle X^{\prime}(t),\operatorname{grad}(F)_{X(t)}\rangle=-\langle Q\operatorname{grad}(F)_{X(t)},\operatorname{grad}(F)_{X(t)}\rangle\leq 0,$
where equality holds if and only if $\operatorname{grad}(F)_{X(t)}=0$ as $Q$ is positive definite, if and only if $X(t)\equiv X_{0}$. In other words, $F$ is a Liapunov function for the dynamical system and it follows from standard methods that $X_{0}$ is asymptotically stable, in the sense that any solution $X(t)$ starting near $X_{0}$ converges to $X_{0}$ as $t\to\infty$ (see e.g. [P, Section 2.9]). On the other hand, if we change to the variable $Y:=Q^{-\tfrac{1}{2}}X$ and consider the new functional $H(Y):=F(Q^{\tfrac{1}{2}}Y)$, then it is easy to check that $X(t)$ is a solution to (14) if and only if $Y(t)=Q^{-\tfrac{1}{2}}X(t)$ is the negative gradient flow of $H$:
$Y^{\prime}(t)=-\operatorname{grad}(H)_{Y(t)}.$
Since $Y_{0}:=Q^{-\tfrac{1}{2}}X_{0}$ is a global minimum and only critical point of $H$ and the function $H$ is real analytic, it is well known that indeed any solution $Y(t)$ converges to $Y_{0}$ as $t\to\infty$.
We therefore obtain that any pluriclosed solution $g(t)$ converges to a bi-invariant metric. More precisely, we have proved the following.
Theorem 5.1.
The pluriclosed solution $g(t)$ on a compact semisimple Lie group $G=G_{1}\times\dots\times G_{s}$ starting at any left and $\operatorname{Ad}(T)$-invariant Hermitian structure $(J,g(0))$, say $g_{\mathfrak{t}}(0)(\mathfrak{t}_{i},\mathfrak{t}_{j})=0$ for all $i\neq j$, $g_{\mathfrak{t}}(0)|_{\mathfrak{t}_{i}}=z_{i}g_{\operatorname{B}}|_{\mathfrak{t}_{i}}$ and $g_{\mathfrak{q}}(0)|_{\mathfrak{q}_{i}}=(z_{i}x_{\alpha})_{\alpha\in\Delta^{+}_{i}}$ as in Theorem 3.7, converges to the bi-invariant metric $g_{b}=z_{1}(-\operatorname{B}_{\mathfrak{g}_{1}})+\dots+z_{s}(-\operatorname{B}_{\mathfrak{g}_{s}})$ as $t\to\infty$.
Recall that in the case when $(G,J)$ is irreducible as a complex manifold, the limit $g_{b}$ is the unique bi-invariant metric (or Bismut flat) up to scaling compatible with $J$.
Example 5.2.
Using (13) and [H, Chapter X, §3] or [K, Appendix C, §1,2], it is straightforward to see that the pluriclosed flow $(x(t),y(t))=(x_{1}(t),x_{2}(t))$ on $G=\mathrm{SU}(3)$ is given by
$\left\{\begin{array}[]{l}x^{\prime}=\frac{2}{x}-\frac{1}{y}+\frac{1}{x+y-1}-2,\\ y^{\prime}=-\frac{1}{x}+\frac{2}{y}+\frac{1}{x+y-1}-2.\end{array}\right$.
Analogously, for $G=\mathrm{SO}(5)$ we obtain that
$\left\{\begin{array}[]{l}x^{\prime}=\frac{2}{x}-\frac{1}{y}+\frac{1}{x+y-1}-2,\\ y^{\prime}=-\frac{1}{x}+\frac{1}{y}+\frac{1}{x+2y-2}-1,\end{array}\right$.
and for $G=\mathrm{G}_{2}$,
$\left\{\begin{array}[]{l}x^{\prime}=\frac{2}{x}-\frac{3}{y}-\frac{1}{x+y-1}+\frac{1}{2x+y-2}+\frac{3}{3x+y-3}-2,\\ y^{\prime}=-\frac{3}{x}+\frac{6}{y}+\frac{3}{x+y-1}-\frac{3}{3x+y-3}+\frac{3}{3x+2y-4}-6.\end{array}\right$.
Note that the lines $x=1$ and $y=1$ are both invariant in the three cases.
6. Appendix: Roots
The following basic properties of root systems can be found in, among many other books, [H, Chapter III, §4,5, Chapter X, §3] and [K, Chapter II, §5].
Given a connected and compact semisimple Lie group $G$ and a maximal torus $T\subset G$, we have that the Lie algebra $\mathfrak{t}$ of $T$ is a maximal abelian subalgebra of $\mathfrak{g}$, the Lie algebra of $G$, and the complexification $\mathfrak{t}^{c}$ is a Cartan subalgebra of $\mathfrak{g}^{c}$. The adjoint action of $\mathfrak{t}^{c}$ on the complex semisimple Lie algebra $\mathfrak{g}^{c}$ determines the root space decomposition:
$\mathfrak{g}^{c}=\mathfrak{t}^{c}\oplus\bigoplus_{\alpha\in\Delta}\mathfrak{g}_{\alpha},$
where $\Delta\subset(\mathfrak{t}^{c})^{*}$ is the root system, i.e., $H,E]=\alpha(H)E$ for all $H\in\mathfrak{t}^{c}$ and $E\in\mathfrak{g}_{\alpha}$. Since the Killing form $\operatorname{B}$ of $\mathfrak{g}^{c}$ is non-degenerate on $\mathfrak{t}^{c}$, for any $\alpha\in\Delta$, there is a unique $H_{\alpha}\in\mathfrak{t}^{c}$ such that $\alpha=\operatorname{B}(H_{\alpha},\cdot)$. There exist vectors $E_{\alpha}$, $\alpha\in\Delta$ such that
$\mathfrak{g}_{\alpha}={\mathbb{C}}E_{\alpha},\qquad\operatorname{B}(E_{\alpha},E_{-\alpha})=1,\qquad H_{\alpha}=[E_{\alpha},E_{-\alpha}],\qquad[E_{\alpha},E_{\beta}]=N_{\alpha,\beta}E_{\alpha+\beta}.$
If $\Delta^{+}$ is an ordering of $\Delta$ (i.e., $\Delta=\Delta^{+}\cup-\Delta^{+}$ (disjoint union) and $(\Delta^{+}+\Delta^{+})\cap\Delta\subset\Delta^{+}$) and $\Pi\subset\Delta^{+}$ is the subset of simple roots (i.e., those which are not the sum of two positive roots), then the set
$\{H_{\alpha}:\alpha\in\Pi\}\cup\{E_{\alpha}:\alpha\in\Delta\}$
is a basis of $\mathfrak{g}^{c}$ called the Chevalley basis and the set
$\{\mathrm{i}H_{\alpha}:\alpha\in\Pi\}\cup\left\{e_{\alpha}:=\tfrac{1}{\sqrt{2}}(E_{\alpha}-E_{-\alpha}),\quad f_{\alpha}:=\tfrac{\mathrm{i}}{\sqrt{2}}(E_{\alpha}+E_{-\alpha})\right\}$
is a basis of $\mathfrak{g}$ called the Cartan-Weyl basis.
Every $\beta\in\Delta^{+}$ can be written in a unique way as $\beta=\sum\limits_{\alpha\in\Pi}n_{\alpha}\alpha$, where $n_{\alpha}\in{\mathbb{Z}}_{\geq 0}$ for any $\alpha\in\Pi$, and there exists a unique maximal root $\alpha_{max}$, i.e., for any $\alpha\in\Delta^{+}$, either $\alpha_{\max}-\alpha\in\Delta^{+}$ or $\alpha_{\max}-\alpha\notin\Delta$. Since $H_{\alpha}\in\mathrm{i}\mathfrak{t}$ for any $\alpha\in\Delta$ and $\operatorname{B}$ is positive definite on $\mathrm{i}\mathfrak{t}$, we can define
$\langle\alpha,\beta\rangle:=\operatorname{B}(H_{\alpha},H_{\beta})=-\operatorname{B}_{\mathfrak{g}}(\mathrm{i}H_{\alpha},\mathrm{i}H_{\beta}),$
where $\operatorname{B}_{\mathfrak{g}}$ is the Killing form of $\mathfrak{g}$.
The vectors $E_{\alpha}$ can be chosen in such a way that the numbers $N_{\alpha,\beta}$ such that $E_{\alpha},E_{\beta}]=N_{\alpha,\beta}E_{\alpha+\beta}$ satisfy that (see [K, Chapter VI, Theorem 6.6])
$N_{-\alpha,-\beta}=-N_{\alpha,\beta}.$
These numbers also satisfy the following properties:
$N_{\beta,\alpha}=-N_{\alpha,\beta},\qquad N_{\alpha,\beta}=N_{\beta,\gamma}=N_{\gamma,\alpha},\quad\mbox{if}\quad\alpha+\beta+\gamma=0,$
and if $\alpha+\beta+\gamma+\delta=0$ and the sum of any pair is nonzero, then
$N_{\alpha,\beta}N_{\gamma,\delta}-N_{\alpha,\gamma}N_{\beta,\delta}+N_{\alpha,\delta}N_{\beta,\gamma}=0.$
Moreover, an explicit formula for $N_{\alpha,\beta}^{2}$ is given as follows. The only multiples of $\alpha\in\Delta$ which are roots are $0,\pm\alpha$, and given $\alpha,\beta\in\Delta$, we have that $\beta+n\alpha\in\Delta$ if and only if $p\leq n\leq q$, where $p\leq 0\leq q$ and
$p+q=-\tfrac{2\langle\alpha,\beta\rangle}{\langle\alpha,\alpha\rangle}\in{\mathbb{Z}},\qquad N_{\alpha,\beta}^{2}=\tfrac{1}{2}q(1-p)\langle\alpha,\alpha\rangle.$
In particular, $N_{\alpha,-\beta}^{2}=N_{\alpha,\beta}^{2}+\langle\alpha,\beta\rangle$, so $N_{\alpha,\beta}^{2}=-\langle\alpha,\beta\rangle$ if $\alpha+\beta\in\Delta$ and $\alpha-\beta\notin\Delta$ (e.g., whenever $\alpha,\beta\in\Pi$).